### v2

# All-weather Multi-Modality Image Fusion: Unified Framework and 100k Benchmark

Abstract
Multi-modality image fusion (MMIF) combines complementary information from different image modalities to provide a more comprehensive and objective interpretation of scenes. However, existing MMIF methods lack the ability to resist different weather interferences in real-world scenes, preventing them from being useful in practical applications such as autonomous driving. To bridge this research gap, we proposed an all-weather MMIF model. Achieving effective multi-tasking in this context is particularly challenging due to the complex and diverse nature of weather conditions. A key obstacle lies in the 'black box' nature of current deep learning architectures, which restricts their multi-tasking capabilities. To overcome this, we decompose the network into two modules: a fusion module and a restoration module. For the fusion module, we introduce a learnable low-rank representation model to decompose images into low-rank and sparse components. This interpretable feature separation allows us to better observe and understand images. For the restoration module, we propose a physically-aware clear feature prediction module based on an atmospheric scattering model that can deduce variations in light transmittance from both scene illumination and reflectance. We also construct a large-scale multi-modality dataset with 100,000 image pairs across rain, haze, and snow conditions, covering various degradation levels and diverse scenes to thoroughly evaluate image fusion methods in adverse weather. Experimental results in both real-world and synthetic scenes show that the proposed algorithm excels in detail recovery and multi-modality feature extraction. The code is available at https://github.com/ixilai/AWFusion.
Index Terms:
Image fusion, adverse weather, multi-modality benchmark, image restoration.I Introduction
Infrared and visible image fusion (IVIF) is performed to integrate the complementary and salient feature information from different modal images to achieve a more comprehensive interpretation of a target scene. By integrating data from different spectral ranges, IVIF contributes to advancements in computer vision tasks [1, 2, 3, 4, 5], including semantic segmentation [6, 7, 8], object detection [9, 10, 11], and unmanned driving [12]. Recently, deep learning (DL) has exhibited remarkable progress within the realm of artificial intelligence owing to its exceptional aptitude for intricate nonlinear approximations and adeptness in extracting salient features. Consequently, several highly efficient algorithms tailored to the IVIF domain have been developed. These algorithms can be broadly categorized into three groups: generative adversarial networks (GAN)-based models [13, 14, 15, 16], autoencoder (AE)-based models [2, 17, 18], and algorithmic expansion models [4, 19].
Each of these methodologies has achieved commendable outcomes in the IVIF field. However, there are three key questions that remain unresolved.
Currently, there is a critical research gap in that IVIF cannot deliver robust scene information under adverse weather conditions. Addressing this challenge requires more than isolated algorithmic solutions for specific weather-related issues, necessitating a unified framework to comprehensively restore intricate scene details affected by adverse weather.
The network design of most current DL-based methods, primarily developed through empirical and experiential approaches, lacks interpretability due to their black-box nature. As a result, identifying an appropriate framework often becomes a time-consuming process.
Generally, DL-based algorithms are driven by data. Existing datasets dedicated to collecting comprehensive imaging data from multiple modalities are complete in terms of scene diversity. However, no large datasets have been developed specifically for research on applying IVIF to scenarios with adverse weather conditions. Even if certain datasets can cover scenes affected by rain and haze, their limited data make it difficult to employ them for training in joint image restoration and fusion.
To address above challenges, we propose an all-weather MMIF algorithm (AWFusion) that integrates both deweathering and fusion functions within a single framework, thereby providing an effective means to achieve robust fusion under adverse weather conditions. Compared to the traditional two-stage strategy of ‘restoration+ fusion’ for addressing adverse weather fusion problems, we propose that an end-to-end framework, which prioritizes modality information interaction before feature recovery, can yield more effective results. The two-stage approach has notable drawbacks: because image restoration and fusion are independent processes, errors introduced during restoration directly impact the fusion outcome. Additionally, the lack of correlation between these tasks means that while efficient restoration may enhance fusion performance, the reverse is not true.
In complex natural scenes, images often contain substantial redundant information, allowing for the inference of the original scene even when the image quality is degraded due to adverse weather conditions. From a network learning perspective, deep neural networks extract features at different levels through layer-by-layer feature extraction operations. Early layers capture low-level features such as edges, textures, and color gradients, while deeper layers learn more complex semantic information and scene structures. However, the loss of low-level features does not entirely compromise high-level scene information. The network can reconstruct the overall scene from remaining valid pixels or local features. Therefore, we argue that it is not necessary to perform image restoration first when handling complex scenes. Second, the task of multi-modality image fusion inherently involves discarding redundant information and retaining complementary data. Unlike image restoration, which aims to recover as much detail as possible without accounting for redundancy, fusion emphasizes retaining only the most significant information. By leveraging complementary multi-modality features, the network can concentrate on these useful details during reconstruction, rather than processing unnecessary pixel data. This not only reduces computational complexity but also enhances the ability of the network to recover the aspects of the scene that best represent its true characteristics. Thus, our objective is to shift the focus of the network from perfect pixel-level recovery to maximizing the representation of key scene information.
The fusion module was constructed to equip the network with the ability to extract multi-modality information, isolate the shared and unique features of each modality, and fuse them based on a unique prior for each feature type. Firstly, we introduce a learned low-rank representation (LLRR) block [4] to decompose the source images into sparse and low-rank components. This feature decoupling approach allows for a more comprehensive consideration of the feature information of the image. Additionally, the low-rank and sparse decompositions reduce the computational complexity of the fusion module. The low-rank component, with its fewer degrees of freedom, represents the global structure with fewer parameters, while the sparse component focuses on a limited number of salient pixels, minimizing unnecessary computational overhead. By leveraging this decomposition, more efficient algorithms can be applied to process large-scale images. For the features integration, a sparse feature prediction module based on an optimization model was proposed to reconstruct the sparse components, and a sparse transformer [20] was introduced to capture the long-range contextual information of an image.
In the restoration module, the clear feature prediction block was developed, guided by the atmospheric scattering model [21, 22]. This module derives illumination and reflectance features from images to predict the distribution of light transmission, utilizing learned convolutional sparse coding (LCSC) [23]. Thus, propelled by both mathematical optimization algorithms and physical models, our proposed model has interpretable and physically-aware properties. It obviates the necessity for model customization to cater to specific weather-related issues, and it enables the simultaneous training of multiple teacher networks under a unified framework for a singular weather restoration task. Subsequently, distillation learning was incorporated into the model to transfer knowledge from these diverse teacher networks to a singular student network, enabling the model to resolve degradation issues stemming from various severe weather phenomena.
Furthermore, we introduce a large-scale benchmark, AWMM-100k, specifically designed for multi-modality image fusion in adverse weather conditions, aiming to advance research in this challenging field. Our dataset consists of 100,000 multi-modality image pairs, comprehensively covering both synthetic and real-world scenes affected by three major weather degradations: rain, haze, and snow. To ensure robust evaluation, the data spans multiple degradation levels for each weather type, capturing the diversity and complexity of natural environments. This benchmark addresses a critical gap in the field by providing a reliable, diverse, and extensive resource for developing and evaluating fusion methods under real-world adverse weather conditions.
The contributions of this study are summarized as follows.
We propose an end-to-end all-weather MMIF framework, effectively fusing and recovering the perturbed multi-modality information. Unlike existing MMIF works, to the best of our knowledge, this is the first work to address MMIF and adverse weather removal simultaneously, defining a new challenging task.
We propose a feature separation-based fusion module that considers the reconstruction of local and global features from sparse and low-rank components.
We propose a physically-aware clear feature prediction block (PFB) that comprehensively accounts for the restoration of clear features, thereby avoiding a time-consuming empirical network design.
Leveraging both existing MMIF datasets and our collected data, we built a dataset (AWMM-100k) dedicated to MMIF research under adverse weather. It containing 100,000 aligned infrared and visible image pairs with labels for object detection, establishing it as a versatile benchmark for a broad spectrum of vision tasks.
II Related Work
II-A DL-based MMIF Methods
DL-based IVIF algorithms [24, 25, 26, 27, 28, 29, 30, 31] demonstrate robust fusion performance in conventional scenes, employing methodologies such as GAN or AE to effectively integrate salient feature information from each modality. With the widespread utilization of IVIF technology, researchers increasingly integrate this technology into downstream vision tasks [3, 32, 33, 34, 35]. Their aim is to enhance the extraction of multi-modality features, augmenting semantic information and leveraging fused images to optimize the performance of tasks such as object detection, semantic segmentation, and salient object detection. Furthermore, there are many joint multi-tasking algorithms designed to cope with real-world applications [36, 37, 38]. In practical applications, images captured by cameras may be slightly misaligned and distorted; some researchers have designed correction networks for fusion specifically for such situations [37]. Xu et al. [38] designed an IVIF model that could fuse unaligned image pairs, effectively reducing the cost of adopting IVIF technology. In some cases, the fixed input and output resolutions of images can limit the development of IVIF-based methods; Li et al. [39] designed an IVIF framework specifically for this problem. In addition, some algorithms apply the IVIF task to degraded environments, such as noise [40], low light [41], and low contrast scenes [42]. However, current research does not effectively address the generalization of IVIF algorithms, whether they are single-task or multi-task, to adverse weather conditions.
III Proposed Method
III-A Overview
As shown in the Fig. 1, AWFusion comprises two main parts: image fusion and image restoration. It is important to emphasize that the proposed algorithm is trained in an end-to-end manner. First, the multi-modality data affected by adverse weather is sent to the image fusion module for feature encoding. To effectively distinguish private and shared features from different modalities and thereby allow the fusion network to overcome the limitations in using a single encoder, a LLRR block is introduced. This block decomposes the source image into a low-rank component, which captures global structural information with high pixel correlation, and a sparse component, which represents high-frequency information with faster pixel variations. This approach enables the targeted handling of specific modal features and enhances fusion outcomes.
In the second part, the multi-modality interaction features obtained from the fusion module are fed into the restoration module to infer the original clear scene. As shown in Fig. 1, the restoration module in the model uses a U-Net backbone. In the backbone network, we stack five sparse transformer blocks (STB) to extract useful features distributed in adverse weather. STB mainly consist of top-k sparse attention (TKSA) and mixed-scale feed-forward network (MSFN). Regarding TKSA, it performs feature aggregation to keep the features in an image sparse, and MSFN extracts multiscale local information from the image [20]. The proposed PFB focuses on refining the spatial information, which is embedded into both sides of the backbone network and enforced before being inputted into subsequent modules, which helps recover the detailed information. Finally, the reconstructed fusion results are output through the restoration module.
Finally, drawing inspiration from Che et al. [43], we introduced a distillation learning approach in which distinct teacher models are initially trained for individual weather-related issues. Subsequently, knowledge is collaboratively transferred from these multiple teacher models to student models to comprehensively address weather-related challenges.
III-B Image Fusion Module
Through the decomposition provided by the LLRR block, our network focuses solely on the fusion of low-rank and sparse features. Low-rank components reflect the global information and main structure of the image, constrained by certain rank features. Given the strong ability of Transformer to capture global context, we employ the STB to fuse low-rank features from different modalities. In contrast, sparse components represent local details such as edges, textures, and high-frequency information. To effectively manage these local fusions, we design the Sparse Feature Block (SFB).
III-B1 SFB
Sparse feature fusion is performed to effectively extract and integrate local features specific to an infrared image () and a visible image (). The fused feature map consists of two parts: the sparse feature for and the sparse feature for . is modeled as follows:
where and are the sparse filters of related to and , respectively, and denotes the number of filters. Subsequently, Eq. 1 can be transformed into the following optimization problem:
Similar to [19], and are solved as follows:
where and are the stack of feature responses and respectively; and are the updates of and at the -th iteration, respectively; and , , and are the learnable convolutional layers. According to LCSC, we can expand the iterations in Eq. 3 and Eq. 4 into a network as shown in the framework in Fig. 2.
III-C Image Restoration Module
The second phase aims to achieve image restoration in the presence of three types of adverse weather conditions (rain, snow, and haze). Based on the atmospheric scattering model, adverse weather conditions such as rain, haze, and snow induce scattering and absorption of light by particles in the atmosphere. Consequently, these atmospheric phenomena lead to comparable image degradation across a spectrum of unfavorable conditions [44]. For instance, during heavy rainfall, the accumulation of distant rain lines and atmospheric water particles creates a ”rain curtain” effect. This visual effect of rainwater accumulation produces a haze like phenomenon in the image background [45]. The above conclusion provide a research basis to realize deweathering in the same framework. Therefore, for the restoration module, an atmospheric scattering model is combined to simulate the process of image degradation, and a PFB is proposed.
III-C1 PFB
The construction of the PFB was guided by the physical mechanism of image degradation and transformation of the atmospheric scattering model. First, utilizing kernel k for feature extraction results in the following revised expression of atmospheric scattering model:
where denotes the convolution operator, and represents the Hadamard product. Subsequently, by introducing the matrix-vector forms of , , , , and , which are denoted as K, J, I, A, and D, respectively, Eq. 5 can be redefined as follows:
According to [21], given the input adverse weather image feature , the clear feature can be computed as follows:
where represents an approximation of the characteristic KA of the atmospheric light , and is a characteristic associated with the transmission map. 1 represents a matrix composed entirely of ones. Since atmospheric light is usually assumed to be homogeneous, as shown in Fig. 1, we use global average pooling to eliminate redundant information within the feature space to predict the atmospheric light feature .
In adverse weather conditions like dense haze, rain, or snow, light undergoes scattering and absorption, causing alterations in both image reflectance and illumination. For example, haze reduces the intensity of light reflected from objects to the camera by scattering it, with this effect being more pronounced at greater distances. Rain and snow also alter the surface reflectivity of objects when they adhere to them, affecting how light is reflected. Additionally, atmospheric particles can cause light to be reflected, refracted, and absorbed, impacting the overall illumination of the scene. Consequently, for the prediction of the feature , by analyzing changes in reflectance and illumination of objects in an image, we can infer changes in light transmission, thus realizing estimation of the image transmission information.
We disassemble into a combination of reflectance and illumination features, which is modeled as follows:
where and are the reflectance and illumination filters of , respectively. Subsequently, we compute the reflectance and illumination feature responses by solving the following optimization problem:
Reflectance Feature Prediction Block (RFB)
The goal of RFB is used to compute the reflectance feature. According to [19], we fix to update by rewriting Eq. 9 as:
Eq. 10 is a standard convolutional sparse coding problem that can be solved using LCSC [23] by rewriting Eq. 10 as:
where is the stack of feature responses , is the update of at the -th iteration, and are the learnable convolutional layers related to the filters , and represents the soft thresholding operator. Similar to SFB, we expand the iterations in Eq. 11 into a network as shown in Fig. 2.
Illumination Feature Prediction Block (IFB)
The goal of IFB is to compute the illuminance feature in a hazy image, the structure of this network is similar to that of DFB. By fixing to update , Eq. 9 can be rewritten as:
By using LCSC to solve the above problem, Eq. 12 can be rewritten as follows:
where is the stack of feature responses , is the update of at the -th iteration, and and are the learnable convolutional layers related to the filters . The network structure of the IFB is shown in Fig. 2.
III-D Loss Function
The loss function integrates two components. The first part focuses on the loss function used to ensure the quality of the fusion results, while the second part addresses the loss function constraining the generation of reflectance and illumination features. The fusion loss includes structural similarity index (SSIM), mean square error (MSE) and color consistency loss. The calculation for SSIM can be expressed as follows:
where and represents two different images. The , where , denotes the mean of image , represents the standard deviation of image , denotes the covariance between images and . The constants and are introduced to prevent the denominator from approaching zero. Thus, the formula for the SSIM loss () is as follow:
where and represent the ground truth of infrared and visible images under adverse weather, respectively. In addition, MSE can be calculated as follows:
where and represent the height and width of the image, respectively. The MSE loss () is as follow:
Colour consistency loss is used to ensure that the fused image maintains consistent colours with the . We transform the image into the YCbCr color space and apply a constraint using the Euclidean distance between the Cb and Cr channels. This can be expressed as:
where represents the transfer function from RGB to the CbCr color space.
To effectively constrain the generation of the reflectance feature and the illuminance feature in Eq. 8, we propose a reconstruction loss inspired by retina theory. According to this theory, a natural image can be viewed as the product of its reflectance and luminance components. Based on this principle, we reconstruct the natural image by multiplying the extracted reflectance and illuminance features, and impose the reconstruction loss using the as the reference. The reconstruction loss can be expressed as follows:
The total loss of AWFusion is denoted as:
Aspect | Description |
Total Image Pairs | 100,000 pairs |
Modalities | Visible (RGB), Infrared (IR) |
Weather Conditions | Rain, Haze, Snow |
Degradation Levels | Heavy, Medium, and Light |
Data Types | Synthetic and Real-world Scenes |
Scenes | Urban, Suburban, Rural, Natural, etc. |
Task Labels | Object Detection |
III-E MMIF Benchmark for Adverse Weather
Current MMIF datasets lack comprehensive scene coverage under adverse weather. Combining existing datasets (RoadScene[28], MSRS[46], M3FD[3], and LLVIP[47]) and our newly collected data in real-world scenes, we have established a benchmark AWMM-100k, which encompasses a diverse range of weather conditions. Each type of weather (rain, haze, and snow) is categorized into heavy, medium, and light levels. Specific descriptions are shown in Table I. The AWMM-100k dataset was collected using a DJI M30T drone, equipped with high-resolution visible and thermal imaging cameras to capture data in complex scenes.
Collecting a large amount of aligned paired multi-modality data and corresponding ground truth in real adverse weather is challenging, leading most current research to rely on synthetic data for training. Effectively degrading clean data and imbuing the model with adaptability to real scenes are pivotal for constructing benchmark. However, the degradation of infrared images in severe weather remains a current research gap. To address this, we analyze data collected during real severe weather to determine the most effective methods for degrading infrared images. Observations of the data in Fig. 3 allow us to draw the following conclusions: 1) While the visible image exhibits degradation due to adverse weather conditions, the corresponding infrared image lacks explicit weather features, showcasing only varying degrees of degradation commensurate with the extent of image disturbance, such as diminished contrast and detail. 2) The depth of the scene affects the extent of light refraction and scattering, resulting in greater blurring in distant scenes compared to closer ones. Based on the above conclusions, we propose an adaptive degradation strategy for infrared images based on scene depth. First, we transpose the image into the Fourier domain and use the depth map as the weights for degradation. Subsequently, the weight map is normalised and multiplied with the amplitude and phase of the image respectively and then inverted to obtain the degraded image. However, distant pixels in the depth map are always represented in black, i.e., the pixel value is . In order to suppress too small weights, we add , and to points with pixel values less than according to the three degrees of light, medium and heavy, respectively.
Degraded visible images generation process is illustrated in Fig. 1. We employ a phenomenological pipeline [48] to introduce haze into source images, while rain and snow images are generated using Photoshop software utilizing masks based approach [49]. In addition, Fig. 4 presents examples of multi-modality data captured in real-world scenarios under adverse weather. Moreover, this benchmark provides labels for object detection task. These attributes render it a versatile resource for investigating a broad spectrum of visual tasks.
IV Experiment
IV-A Experimental Configurations
We performed fusion experiments on the AWMM-100k dataset under adverse weather (rain, haze, and snow). The experiment included ten state-of-the-art (SOTA) IVIF algorithms (U2Fusion[2], LRRNet[4], MURF[38], TIM[35], CoCoNet[50], TarDal[3], IGNet[32], DRMF[51], MUFusion[52], CrossFuse[53]). We chose nine metrics to quantitatively evaluate different fusion results: Gradient-Based Fusion Performance (), Image Fusion Metric-Based on a Multiscale Scheme (), Piella’s Metric (), Chen-Blum Metric (), Chen-Varshney Metric () [1, 54], [55], Structural Similarity Measurement (), the Sum of the Correlations of Differences ()[56] and Entropy (). For the synthesized data, the comparison method uses clean source images, while the proposed algorithm performs fusion using images affected by three different adverse weather conditions. For real scene data, both the proposed algorithm and the comparison method use the same source images. This comparison setup allows for a more effective demonstration of the performance of the proposed algorithm.
IV-B Training Details
For the training process, three specialized teacher networks and one student network were trained under different weather conditions. For different adverse weather conditions, we selected 1,000 pairs of multi-modality images from the AWMM-100k dataset as the training set for the teacher network. Each image was randomly cropped to a size of pixels to form a batch. The initial learning rate for the Adam optimizer was set to , and the model training was conducted over 300 iterations. During the knowledge transfer phase, the teacher network remained fixed while the student network was trained for 200 epochs on a mixed training set comprising three weather types. Furthermore, we employed two LLRR blocks, while the remaining parameters were set according to the original paper. The number of LCSC blocks iterated in the RFB, IFB, and SFB models is four, with 24 filters per block. The size of the convolutional kernel used in each convolutional layer is nine. All experiments were carried out using the PyTorch framework on a system equipped with two NVIDIA 4090 GPUs.
IV-C Qualitative Comparison
IV-C1 Comparison on synthetic data
Figs. 5, 6 and 7 present the fusion results of different algorithms on the synthetic dataset. Notably, only the proposed algorithm performs fusion under adverse weather conditions, while the comparison methods are fused in ideal scenes. Additionally, we include pseudo-color images of the disturbed visible images and the fusion results from the student model to further demonstrate the effectiveness of the proposed algorithm in mitigating severe weather interference. AWFusion-T and AWFusion-S represent the teacher and student models of the proposed algorithm, respectively. As shown in Figs. 5, 6 and 7, all comparison methods achieve good fusion results in ideal scenes. Four methods—U2Fusion, LRRNet, MURF, and IGNet—produce fusion results with low contrast. In contrast, CoCoNet and MUFusion enhance the source image features, leading to some redundancy and loss of detail in the thermal radiation information in the fusion results. In daytime scenes, the TIM, TarDal, and DRMF methods effectively retain the texture information from the visible images. However, in nighttime scenes, the TarDal method exhibits a loss of detail. Additionally, both MUFusion and CrossFuse show poor color retention of the sky, introducing excessive infrared information into the region, which is not compatible with the human visual system.
Under severe weather interference, the challenge not only arises from weather-related artifacts on the visible images but also from the low contrast of infrared images and the loss of thermal radiation information. By examining the fusion results of the proposed algorithm in adverse weather conditions, we observe that it performs effective, high-quality fusion by removing interfering information from the visible image and enhancing weak targets in the scene. Additionally, in the red zoomed-in areas, the algorithm not only recovers a clean scene but also facilitates multi-modality information interactions, integrating salient details from different modalities. This results in superior performance compared to most of the comparison methods, particularly in terms of overall contrast and detail preservation. Furthermore, the pseudo-colored images show that the proposed algorithm effectively removes interference without smoothing local texture information, which is crucial for demonstrating the restoration performance of algorithm. Since the student model exhibits stronger generalization ability compared to the teacher model, some performance is sacrificed. From the experiments, we observe that the first noticeable effect is the weakening of color information, followed by the retention of infrared information in the sky region. However, this does not affect the semantic content of the overall image, and the student model remains close to the teacher model in terms of overall contrast.
IV-C2 Comparison on real-world data
Fig. 9, we present the fusion results of all methods in a real haze scene. Additionally, we evaluate the fusion performance of different methods using two downstream tasks: object detection and depth estimation. We used YOLOv7 [10] as a detector and MiDas [57] as a depth estimator. Unlike synthetic data, the input images used to comparison methods and the proposed algorithm are the same, as there is no corresponding ground truth for real scene data. Furthermore, we retrained three high-performing dehaze algorithms (Dehamer[58], DehazeFormer[59], UCL-Dehaze[60]) using images from the AWMM-100K to optimize the visible images for input into the fusion comparison methods. However, as seen in Fig. 8, all dehaze methods resulted in varying degrees of color distortion, suggesting that feeding such distorted inputs into the different fusion methods could lead to unfair comparisons. Therefore, we provided all fusion methods with uniform input images.
As shown in Fig. 9, in the first set of fusion comparison tests, CoCoNet, TarDal, and MUFusion effectively retain the significant thermal radiation information from the infrared map. These three methods, which tend to enhance infrared data to achieve higher contrast, can address some challenges in real haze scenes. However, this approach excessively masks details from the visible image, preventing the recovery of clear scene details. Additionally, for pedestrians submerged in haze, DRMF and CrossFuse lose target information from the infrared image, incorrectly prioritizing the blurred pedestrian details from the visible image. In contrast, the proposed algorithm strikes an effective balance between the complementary information from different modalities, maintaining good overall scene contrast while also recovering some detail in unknown real scenes. By examining the accuracies of different methods in the object detection task, it is clear that only the AWFusion-S results successfully detect all target information in the scene. While the CoCoNet method enhances target information, this also introduces detection errors, suggesting that blind enhancement is not an effective solution to the problem of degraded fusion. The proposed algorithm effectively recovers scene details and integrates complementary multi-modality information, achieving the highest detection accuracy across all objects. For the depth estimation task, the results visualize each ability of method to extract overall scene information. While all methods preserved depth in the foreground, many, such as U2Fusion, LRRNet, CoCoNet, and MUFusion, struggled with accurate depth estimation for distant pedestrians obscured by haze. Additionally, haze affected the performance of TIM, IGNet, and DRMF, as these methods failed to maintain contrast between the roadside fence and its surroundings, leading to edge blurring in the depth map.
In summary, the proposed algorithm outperforms current SOTA methods in both fusion performance and semantic information retention when addressing real-world scene challenges. It demonstrates excellent performance in downstream tasks, such as object detection and depth estimation, indicating its high practical application value in real-world.
Source Method Pub. ↑ ↑ ↑ ↑ ↓ ↑ ↑ ↑ ↑ U2Fusion[2] TPAMI 2020 0.2548 0.4524 0.7769 0.4565 901.4884 0.2985 0.3631 1.0276 4.4955 LRRNet[4] TPAMI 2023 0.3194 0.5046 0.7003 0.3593 682.5369 0.4250 0.1708 0.7897 5.8903 MURF[38] TPAMI 2023 0.2570 0.3839 0.5732 0.3218 1523.9926 0.2285 0.2699 0.9151 4.5988 TIM[35] TPAMI 2024 0.2057 0.3498 0.4630 0.3844 827.1437 0.3035 0.1585 0.9651 6.1997 CoCoNet[50] IJCV 2023 0.1526 0.2232 0.4021 0.4530 1099.7753 0.2129 0.1756 1.3002 7.6143 TarDal[3] CVPR 2022 0.3563 0.3859 0.5739 0.2710 653.6011 0.3826 0.3053 1.5556 6.2024 IGNet[32] MM 2023 0.3529 0.4921 0.7232 0.4570 919.9543 0.4420 0.3515 1.5791 6.0812 DRMF[51] MM 2024 0.3108 0.4170 0.5844 0.5292 929.2306 0.3555 0.2743 0.8937 6.2728 MUFusion[52] InfFus 2023 0.3438 0.4696 0.7860 0.5102 1039.5903 0.4199 0.3959 1.4061 5.6906 GT CrossFuse[53] InfFus 2024 0.4057 0.5718 0.7132 0.4655 1840.1690 0.4648 0.3832 1.0081 5.9859 AWFusion-T 0.3582 0.6855 0.8444 0.4613 464.7867 0.4839 0.3693 1.6311 6.3160 Rain AWFusion-S - 0.3334 0.6428 0.8335 0.4691 476.3436 0.4618 0.3528 1.6010 6.2777 U2Fusion[2] TPAMI 2020 0.3654 0.4275 0.7236 0.5051 840.3653 0.4086 0.3911 1.1512 5.7205 LRRNet[4] TPAMI 2023 0.4154 0.4775 0.7646 0.4455 463.4354 0.5218 0.3046 0.8130 6.6553 MURF[38] TPAMI 2023 0.2589 0.3324 0.6395 0.4055 1305.4273 0.2277 0.3124 0.9431 5.3601 TIM[35] TPAMI 2024 0.3020 0.3983 0.6499 0.4135 507.1376 0.4078 0.2833 0.8978 7.1141 CoCoNet[50] IJCV 2023 0.2664 0.2446 0.6157 0.4647 556.1065 0.3547 0.2987 1.4288 7.6564 TarDal[3] CVPR 2022 0.4109 0.3832 0.7308 0.3710 337.3376 0.4366 0.4009 1.4207 6.8040 IGNet[32] MM 2023 0.3271 0.4117 0.5225 0.4234 1293.6847 0.4110 0.2683 1.4465 6.2330 DRMF[51] MM 2024 0.5203 0.6138 0.7530 0.4463 477.8087 0.5026 0.4182 1.0414 6.5732 MUFusion[52] InfFus 2023 0.3629 0.4200 0.7236 0.4908 2079.8455 0.4451 0.3723 1.2420 6.5892 GT CrossFuse[53] InfFus 2024 0.4542 0.5423 0.7301 0.4559 2474.6507 0.4847 0.3850 0.9437 6.5648 AWFusion-T 0.4497 0.5889 0.7949 0.4838 374.9215 0.5585 0.4078 1.2209 6.9558 Haze AWFusion-S - 0.4283 0.5471 0.7738 0.4610 462.8585 0.5322 0.3901 1.1719 6.9469 U2Fusion[2] TPAMI 2020 0.3035 0.4330 0.7349 0.4547 1031.8395 0.3606 0.3666 1.0589 5.1796 LRRNet[4] TPAMI 2023 0.3574 0.4651 0.7154 0.3737 661.1762 0.4719 0.2361 0.8301 6.1878 MURF[38] TPAMI 2023 0.2440 0.3547 0.5581 0.3482 1613.7692 0.2180 0.2676 0.8499 5.0984 TIM[35] TPAMI 2024 0.2433 0.3626 0.5181 0.3739 820.5891 0.3465 0.2105 0.8977 6.5951 CoCoNet[50] IJCV 2023 0.2420 0.2580 0.5238 0.4512 828.9278 0.3277 0.2593 1.4450 7.5726 TarDal[3] CVPR 2022 0.3511 0.3718 0.6238 0.2903 553.2230 0.4171 0.3453 1.4684 6.4487 IGNet[32] MM 2023 0.3268 0.4577 0.5938 0.4324 1141.8494 0.4356 0.2986 1.5142 6.0336 DRMF[51] MM 2024 0.3948 0.4684 0.6747 0.5055 729.0171 0.4649 0.3417 1.2065 6.2139 MUFusion[52] InfFus 2023 0.3543 0.4336 0.7342 0.4953 1699.4374 0.4372 0.3787 1.2984 6.1981 GT CrossFuse[53] InfFus 2024 0.4090 0.5187 0.6995 0.4567 2421.3634 0.4507 0.3724 0.9919 6.1354 AWFusion-T 0.3580 0.5591 0.8179 0.4827 246.8077 0.4985 0.3780 1.5700 6.6173 Snow AWFusion-S - 0.3552 0.5535 0.8119 0.4798 270.1595 0.4937 0.3734 1.4937 6.6531
IV-D Quantitative Comparison
IV-D1 Fusion Performance Comparison
We selected 50 pairs of infrared and visible source images from the AWMM 100k dataset for each of the three types of adverse weather: rain, haze, and snow, as the test set. The ground truth corresponding to the degraded images was used as input for all comparison methods. Table II presents the results of all methods across nine objective metrics, with the first column indicating the source image types used by each method. The top three scores for each metric are highlighted. Notably, Table II shows that the proposed method, even when performing fusion under adverse weather conditions, outperforms comparison methods operating under ideal conditions. In rain, haze, and snow, the proposed teacher model ranks in the top three for eight or more metrics, while the student model ranks in the top three for , , and metrics, respectively.
By specifically analyzing the scores of each metric, we can draw the following conclusions:
(1) In the image feature-based metrics and , which are primarily computed using the Sobel edge operator and two-level Haar wavelet, respectively, these metrics assess the edge information in the fusion results. The proposed algorithm achieves high scores in both, indicating that AWFusion preserves the edge structure while effectively removing degradation (such as haze) that significantly obscures edges.
(2) Analyzing the and metrics, it is worth noting that AWFusion-S performs slightly worse in rain and haze scenes compared to snow on the metric. This discrepancy arises because rain and haze are continuously distributed degradations, while snow is more localized. As a result, snow is slightly easier to process than rain and haze. The proposed algorithm does not rank among the top three in the metric for haze scenes. This metric evaluates image quality based on texture, local structural changes, and noise sensitivity, suggesting that the performance of algorithm in extracting local texture under haze is somewhat weaker than in other conditions.
(3) For the metrics and , which are inspired by human perception, the proposed method performs exceptionally well. This aligns with the results shown in Figs. 5, 6 and 7 and Fig. 9, where the method effectively integrates multi-modality information while eliminating interference, ensuring that the fusion results are not distorted by adverse weather or enhanced in ways that conflict with human visual perception.
In summary, the performance of the proposed algorithm in quantitative comparisons further validates its practical value in handling severe weather conditions, demonstrating excellent fusion capabilities.
Source | Method | Pub. | Background | Car | Person | Bike | Curve | Car Stop | Cuardrail | Color cone | Bump | mIoU |
U2Fusion[2] | TPAMI 2020 | 98.39 | 92.88 | 73.25 | 73.03 | 52.82 | 73.75 | 86.2 | 40.31 | 83.2 | 74.87 | |
LRRNet[4] | TPAMI 2023 | 98.51 | 93.18 | 69.62 | 73.89 | 65.18 | 77.95 | 84.59 | 35.95 | 88.96 | 76.43 | |
MURF[38] | TPAMI 2023 | 99.13 | 95.56 | 83.54 | 86.65 | 76.27 | 88.21 | 93.48 | 76.25 | 90.59 | 87.74 | |
TIM[35] | TPAMI 2024 | 98.22 | 91.17 | 67.94 | 69.4 | 59.53 | 67.74 | 83.42 | 30.59 | 85.59 | 72.62 | |
CoCoNet[50] | IJCV 2023 | 98.3 | 92.03 | 69.31 | 70.66 | 57.72 | 73.62 | 87.49 | 29.44 | 84.9 | 73.72 | |
TarDal[3] | CVPR 2022 | 98.39 | 92.29 | 70 | 68.47 | 62.07 | 74.59 | 87.9 | 35.65 | 85.42 | 75.01 | |
IGNet[32] | MM 2023 | 99.12 | 95.61 | 83.09 | 87.22 | 74.97 | 87.96 | 93.1 | 76.13 | 90.9 | 87.6 | |
DRMF[51] | MM 2024 | 99.16 | 95.14 | 82.36 | 88.2 | 84.27 | 90.68 | 91.67 | 80.35 | 92.08 | 89.32 | |
MUFusion[52] | InfFus 2023 | 99.23 | 95.17 | 83.84 | 88.43 | 84.1 | 88.9 | 92.94 | 80.7 | 86.12 | 88.82 | |
GT | CrossFuse[53] | InfFus 2024 | 98.04 | 90.3 | 67.43 | 68.51 | 58.58 | 71.81 | 68.13 | 29.92 | 68.7 | 69.04 |
Rain | 99.32 | 96.49 | 84.77 | 88.76 | 86.35 | 90.45 | 93.91 | 78.45 | 91.82 | 90.04 | ||
Haze | 99.19 | 95.69 | 81.51 | 81.94 | 85.21 | 89.5 | 91.57 | 72.73 | 93.57 | 87.88 | ||
Snow | AWFusion-S | - | 99.25 | 95.92 | 81.24 | 86.4 | 86.77 | 90.36 | 93.74 | 78.04 | 92.32 | 89.34 |
Source | Method | Pub. | People | Car | Bus | Lamp | Motorcycle | Truck | mAP@0.5 | mAP@[0.5:0.95] |
U2Fusion[2] | TPAMI 2020 | 0.816 | 0.903 | 0.901 | 0.716 | 0.711 | 0.801 | 0.808 | 0.512 | |
LRRNet[4] | TPAMI 2023 | 0.785 | 0.909 | 0.912 | 0.764 | 0.721 | 0.808 | 0.817 | 0.484 | |
MURF[38] | TPAMI 2023 | 0.815 | 0.885 | 0.881 | 0.711 | 0.675 | 0.8 | 0.794 | 0.498 | |
TIM[35] | TPAMI 2024 | 0.731 | 0.898 | 0.878 | 0.657 | 0.656 | 0.769 | 0.765 | 0.48 | |
CoCoNet[50] | IJCV 2023 | 0.791 | 0.889 | 0.889 | 0.696 | 0.641 | 0.762 | 0.778 | 0.49 | |
TarDal[3] | CVPR 2022 | 0.792 | 0.877 | 0.857 | 0.585 | 0.62 | 0.713 | 0.741 | 0.462 | |
IGNet[32] | MM 2023 | 0.793 | 0.859 | 0.836 | 0.532 | 0.554 | 0.721 | 0.716 | 0.439 | |
DRMF[51] | MM 2024 | 0.699 | 0.897 | 0.896 | 0.787 | 0.697 | 0.784 | 0.793 | 0.494 | |
MUFusion[52] | InfFus 2023 | 0.803 | 0.906 | 0.908 | 0.734 | 0.694 | 0.802 | 0.808 | 0.507 | |
CrossFuse[53] | InfFus 2024 | 0.692 | 0.891 | 0.871 | 0.671 | 0.674 | 0.771 | 0.762 | 0.473 | |
GT | AWFusion-S | - | 0.791 | 0.922 | 0.93 | 0.804 | 0.726 | 0.821 | 0.832 | 0.534 |
IV-D2 Semantic Segmentation Performance Comparison
The proposed all-weather MMIF algorithms are designed to support real-world applications, such as autonomous driving, surveillance, and UAV navigation. In these scenarios, the goal is not just to generate visually enhanced images but to enable effective decision-making and analysis (e.g., object detection, semantic segmentation) from the fused images. To comprehensively evaluate the fusion performance under three types of severe weather, we introduced a downstream task. In Table III, we compare the performance of various algorithms using BANet [61] as the segmentation network. Unlike comparison methods, which receive clean source images as input, the proposed algorithm performs fusion under severe weather conditions. As shown in Table III, the proposed algorithm ranks first in mean intersection-over-union (mIoU) for rain scenes and second for snow scenes. Although the mIoU score for the haze scenario is not in the top three, it still secures fifth place. These experiments demonstrate that the proposed algorithm retains high-level semantic information in scenes with interference and accurately recognizes key targets such as roads, pedestrians, and vehicles.
IV-D3 Object Detection Performance Comparison
In real-world applications, systems must adapt to varying weather and environmental conditions. Although our method is tailored for severe weather, it also delivers high-quality fusion results in normal conditions, such as clear, haze-free, and rain-free environments. This demonstrates not only its effectiveness in harsh weather but also its strong generalization ability in clean settings. In Table IV, we evaluated the object detection accuracies of all methods on the M3FD dataset using YOLOv7 [10]. As shown in Table IV, the proposed algorithm achieves the highest detection accuracy across all five categories, with both mAP@0.5 and mAP@[0.5:0.95] also obtaining the top scores. These results confirm that the proposed algorithm retains sufficient information and detail for accurate localization and identification in object detection. Moreover, it enhances detection accuracy by preserving and leveraging complementary information.
IV-E Extension to Multiple Adverse Weather Disturbances
Compound degradation increases scene complexity, as rain, haze, and snow exhibit distinct physical characteristics: rain and snow cause localized occlusion, while haze reduces overall contrast. When combined, these effects create more challenging degradation. Although specific patterns can address individual disturbances, the model must exhibit strong adaptability and generalization to handle multiple disturbances simultaneously. To evaluate the generalization ability of the proposed algorithm under severe weather conditions, we conducted extended experiments with composite degradation. We used DT [62] as a preprocessing model to remove interference from visible images. In the comparison method, fusion is performed on the preprocessed image, while the proposed algorithm directly processes the fusion result on the original source image. As shown in Fig. 10, when two disturbances coexist in the source image, DT can only address one at a time, despite its effectiveness handle individual interferences. The residual degradation is carried over to the fusion result, leading to suboptimal outcomes. In contrast, the proposed method reconstructs high-fidelity scene details under complex weather conditions, preserving key features while avoiding issues like insufficient brightness or information loss.
IV-F Ablation Experiment
We performed an ablation analysis on the six modules to verify their individual roles and effectiveness. The modules were grouped into four experimental sets, with the results visualized in Fig. 11 and quantitatively compared in Table V.
IV-F1 Effectiveness Analysis of the Fusion Module
The image fusion module plays a critical role in the end-to-end framework by enabling multi-modality feature interaction. To evaluate its effectiveness, we performed an ablation analysis. As illustrated in the first row of Fig. 11, removing the fusion module disrupts this interaction, causing salient infrared features to be occluded by blurred visible details, as seen in the green zoomed-in area. Additionally, the absence of feature interaction introduces low-contrast IR information into the fusion output, making it difficult to reconstruct color information and resulting in slight color distortion, highlighted in the red zoomed-in area.
IV-F2 Effectiveness Analysis of the PFB
The PFB module, built on the atmospheric scattering model, effectively mitigates weather-induced interference and reconstructs clear scene features. We performed an ablation analysis. As shown in the second row of Fig. 11, removing the PFB module leaves residual degradation in the fusion result. In texture-rich regions, such as leaves, the network struggles to distinguish between degradation artifacts and actual textures, as highlighted in the red zoomed-in area, leading to detail loss and blurred edges.
w/o PFB | 0.5439 | 0.7917 | 579.3986 | 0.3048 | 6.2609 |
w/o Fusion | 0.4989 | 0.7953 | 577.4851 | 0.2165 | 6.2930 |
w/o IFB and RFB | 0.4646 | 0.8096 | 501.8484 | 0.2923 | 5.9288 |
w/o STB and SFB | 0.5454 | 0.7979 | 536.2539 | 0.3904 | 6.7516 |
AWFusion | 0.6855 | 0.8444 | 464.7867 | 0.3993 | 6.3160 |
IV-F3 Effectiveness Analysis of the IFB and RFB
To improve the ability of network to estimate light transmission, we designed the IFB and RFB modules based on retina theory and constrained feature generation by reconstructing the loss. As shown in the third row of Fig. 11, removing these modules results in reduced lighting in the fusion output, likely due to inaccurate transmission map estimation and insufficient luminance compensation. This makes the absence of light recovery more noticeable. Additionally, Fig. 4 shows a significant drop in the , , and metrics.
IV-F4 Effectiveness Analysis of the Fusion Strategies
To leverage the low-rank and sparse features decomposed by the LLRR block, we designed the SFB to fuse sparse components and the STB to integrate low-rank components through context extraction. An ablation study was performed by removing the SFB and STB, replacing them with a maximum-value fusion strategy. As more feature channels are extracted from the source image, the simple fusion strategy impedes effective cross-modal interaction, restricting the extraction of salient and complementary features. This causes pixel redundancy in the fused output, reflected in the excessively high metric in Table V. The enlarged region in Fig. 11 further highlights increased noise in ground, complicating object recognition.
V Conclusion
In this study, we propose an all-weather MMIF method, focusing on the relationship between multi-modality image fusion and image restoration. Our framework shifts the network’s objective from pixel-perfect recovery to maximizing the representation of key scene information. The end-to-end design comprises two main modules: a fusion module and a restoration module. The fusion module decouples features from multiple modalities into low-rank and sparse components, which are merged interactively using specialized fusion strategies. Guided by a physical model, the restoration module restores clear details from blur pixels, ensuring high-quality fusion results. Additionally, we introduce a large-scale benchmark dataset, AWMM-100k, tailored for multi-modality image processing under adverse weather conditions. Experimental results confirm that the AWFusion surpasses existing methods in both adverse and normal conditions, with outstanding performance in downstream tasks, demonstrating its strong practical value.

### v3

@@
 Abstract
-Multi-modality image fusion (MMIF) combines complementary information from different image modalities to provide a more comprehensive and objective interpretation of scenes.
-However, existing MMIF methods lack the ability to resist different weather interferences in real-world scenes, preventing them from being useful in practical applications such as autonomous driving.
-To bridge this research gap, we proposed an all-weather MMIF model.
-Achieving effective multi-tasking in this context is particularly challenging due to the complex and diverse nature of weather conditions.
-A key obstacle lies in the 'black box' nature of current deep learning architectures, which restricts their multi-tasking capabilities.
-To overcome this, we decompose the network into two modules: a fusion module and a restoration module.
-For the fusion module, we introduce a learnable low-rank representation model to decompose images into low-rank and sparse components.
-This interpretable feature separation allows us to better observe and understand images.
-For the restoration module, we propose a physically-aware clear feature prediction module based on an atmospheric scattering model that can deduce variations in light transmittance from both scene illumination and reflectance.
-We also construct a large-scale multi-modality dataset with 100,000 image pairs across rain, haze, and snow conditions, covering various degradation levels and diverse scenes to thoroughly evaluate image fusion methods in adverse weather.
-Experimental results in both real-world and synthetic scenes show that the proposed algorithm excels in detail recovery and multi-modality feature extraction.
-The code is available at https://github.com/ixilai/AWFusion.
-Index Terms:
-Image fusion, adverse weather, multi-modality benchmark, image restoration.I Introduction
+Multi-modality image fusion (MMIF) combines complementary information from different image modalities to provide a comprehensive and objective interpretation of scenes.
+However, existing fusion methods cannot resist different weather interferences in real-world scenes, limiting their practical applicability.
+To bridge this gap, we propose an end-to-end, unified all-weather MMIF model.
+Rather than focusing solely on pixel-level recovery, our method emphasizes maximizing the representation of key scene information through joint feature fusion and restoration.
+Specifically, we first decompose images into low-rank and sparse components, enabling effective feature separation for enhanced multi-modality perception.
+During feature recovery, we introduce a physically-aware clear feature prediction module, inferring variations in light transmission via illumination and reflectance.
+Clear features generated by the network are used to enhance salient information representation.
+We also construct a large-scale MMIF dataset with 100,000 image pairs comprehensively across rain, haze, and snow conditions, as well as covering various degradation levels and diverse scenes.
+Experimental results in both real-world and synthetic scenes demonstrate that the proposed method excels in image fusion and downstream tasks such as object detection, semantic segmentation, and depth estimation.
+The source code is available at https://github.com/ixilai/AWFusion.
+keywords:
+Image fusion, Adverse weather, Multi-modality benchmark, Image restoration1 Introduction
 Infrared and visible image fusion (IVIF) is performed to integrate the complementary and salient feature information from different modal images to achieve a more comprehensive interpretation of a target scene.
-By integrating data from different spectral ranges, IVIF contributes to advancements in computer vision tasks [1, 2, 3, 4, 5], including semantic segmentation [6, 7, 8], object detection [9, 10, 11], and unmanned driving [12].
-Recently, deep learning (DL) has exhibited remarkable progress within the realm of artificial intelligence owing to its exceptional aptitude for intricate nonlinear approximations and adeptness in extracting salient features.
-Consequently, several highly efficient algorithms tailored to the IVIF domain have been developed.
-These algorithms can be broadly categorized into three groups: generative adversarial networks (GAN)-based models [13, 14, 15, 16], autoencoder (AE)-based models [2, 17, 18], and algorithmic expansion models [4, 19].
-Each of these methodologies has achieved commendable outcomes in the IVIF field. However, there are three key questions that remain unresolved.
-Currently, there is a critical research gap in that IVIF cannot deliver robust scene information under adverse weather conditions.
+By integrating data from different spectral ranges, IVIF contributes to advancements in computer vision tasks [32, 17, 38, 33, 28], including semantic segmentation [24, 3], object detection [54, 55], and unmanned driving [62].
+Recently, deep learning (DL) has exhibited remarkable progress within the realm of artificial intelligence [65], owing to its exceptional aptitude for intricate nonlinear approximations and adeptness in extracting salient features.
+Consequently, several algorithms tailored to the IVIF have been developed.
+These algorithms can be broadly categorized into four groups: generative adversarial networks (GAN)-based models [18, 67], diffusion-based models [75, 69], autoencoder (AE)-based models [73, 53], and algorithmic expansion models [23, 10].
+However, most of these methods are designed for ideal imaging conditions.
+With the increasing need for robust perception in real-world applications, IVIF is expected to deliver reliable scene representation under adverse weather, where visibility is reduced and modality information becomes severely imbalanced.
+Although current DL-based fusion methods [60, 2, 7, 30] achieve promising results in clear-weather scenarios, their generalization and robustness remain limited in rainy, hazy, or snowy environments.
+Therefore, three major challenges need to be addressed:
+Currently, there is a critical research gap in that IVIF cannot deliver scene information under adverse weather.
 Addressing this challenge requires more than isolated algorithmic solutions for specific weather-related issues, necessitating a unified framework to comprehensively restore intricate scene details affected by adverse weather.
@@
 However, no large datasets have been developed specifically for research on applying IVIF to scenarios with adverse weather conditions.
-Even if certain datasets can cover scenes affected by rain and haze, their limited data make it difficult to employ them for training in joint image restoration and fusion.
-To address above challenges, we propose an all-weather MMIF algorithm (AWFusion) that integrates both deweathering and fusion functions within a single framework, thereby providing an effective means to achieve robust fusion under adverse weather conditions.
-Compared to the traditional two-stage strategy of ‘restoration+ fusion’ for addressing adverse weather fusion problems, we propose that an end-to-end framework, which prioritizes modality information interaction before feature recovery, can yield more effective results.
-The two-stage approach has notable drawbacks: because image restoration and fusion are independent processes, errors introduced during restoration directly impact the fusion outcome.
-Additionally, the lack of correlation between these tasks means that while efficient restoration may enhance fusion performance, the reverse is not true.
-In complex natural scenes, images often contain substantial redundant information, allowing for the inference of the original scene even when the image quality is degraded due to adverse weather conditions.
+In summary, it is essential to build a large-scale multi-modality image dataset for specifically tailored for adverse weather scenarios, and to develop an effective unified “Restoration + Fusion” framework, enabling more reliable IVIF performance in challenging environments.
+1.1 Our Contributions
+To address above challenges, we propose an all-weather MMIF algorithm (AWFusion) that integrates both deweathering and fusion functions within a single framework, thereby providing an effective means to achieve robust fusion under adverse weather.
+To clearly present the design principles of the proposed algorithm, we describe it from the following three aspects.
+(1) Feasibility Analysis of Unified Framework: We argue that to construct a effective unified framework, complementary information should first be integrated without redundancy, followed by restoring the overall clear features.
+The main reasons are as follow: (1) In natural scenes, images contain sufficient redundancy, allowing the original content to be inferred even under adverse weather degradation.
 From a network learning perspective, deep neural networks extract features at different levels through layer-by-layer feature extraction operations.
-Early layers capture low-level features such as edges, textures, and color gradients, while deeper layers learn more complex semantic information and scene structures.
+Early layers capture low-level cues like edges and textures, while deeper layers encode complex semantics and scene structures.
 However, the loss of low-level features does not entirely compromise high-level scene information.
 The network can reconstruct the overall scene from remaining valid pixels or local features.
-Therefore, we argue that it is not necessary to perform image restoration first when handling complex scenes.
-Second, the task of multi-modality image fusion inherently involves discarding redundant information and retaining complementary data.
-Unlike image restoration, which aims to recover as much detail as possible without accounting for redundancy, fusion emphasizes retaining only the most significant information.
-By leveraging complementary multi-modality features, the network can concentrate on these useful details during reconstruction, rather than processing unnecessary pixel data.
+Therefore, it is not necessary to perform image restoration first when handling complex scenes.
+(2) The task of IVIF inherently involves discarding redundant information and retaining complementary data.
+Unlike image restoration, which aims to recover all details, image fusion focuses on preserving the most significant information.
+By leveraging complementary multi-modality features, the network emphasizes useful details while ignoring redundant pixels.
 This not only reduces computational complexity but also enhances the ability of the network to recover the aspects of the scene that best represent its true characteristics.
 Thus, our objective is to shift the focus of the network from perfect pixel-level recovery to maximizing the representation of key scene information.
-The fusion module was constructed to equip the network with the ability to extract multi-modality information, isolate the shared and unique features of each modality, and fuse them based on a unique prior for each feature type.
-Firstly, we introduce a learned low-rank representation (LLRR) block [4] to decompose the source images into sparse and low-rank components.
+The unified framework is consisting of two main components: a fusion module and a restoration module.
+(2) Implementation of Unified Framework: In the fusion module, we introduce a learned low-rank representation (LLRR) block [23] to decompose the source images into sparse and low-rank components firstly.
 This feature decoupling approach allows for a more comprehensive consideration of the feature information of the image.
-Additionally, the low-rank and sparse decompositions reduce the computational complexity of the fusion module.
-The low-rank component, with its fewer degrees of freedom, represents the global structure with fewer parameters, while the sparse component focuses on a limited number of salient pixels, minimizing unnecessary computational overhead.
-By leveraging this decomposition, more efficient algorithms can be applied to process large-scale images.
-For the features integration, a sparse feature prediction module based on an optimization model was proposed to reconstruct the sparse components, and a sparse transformer [20] was introduced to capture the long-range contextual information of an image.
-In the restoration module, the clear feature prediction block was developed, guided by the atmospheric scattering model [21, 22].
-This module derives illumination and reflectance features from images to predict the distribution of light transmission, utilizing learned convolutional sparse coding (LCSC) [23].
-Thus, propelled by both mathematical optimization algorithms and physical models, our proposed model has interpretable and physically-aware properties.
-It obviates the necessity for model customization to cater to specific weather-related issues, and it enables the simultaneous training of multiple teacher networks under a unified framework for a singular weather restoration task.
-Subsequently, distillation learning was incorporated into the model to transfer knowledge from these diverse teacher networks to a singular student network, enabling the model to resolve degradation issues stemming from various severe weather phenomena.
-Furthermore, we introduce a large-scale benchmark, AWMM-100k, specifically designed for multi-modality image fusion in adverse weather conditions, aiming to advance research in this challenging field.
+Additionally, low-rank and sparse decomposition reduces the computational cost of fusion module: the low-rank component can capture global structure with fewer parameters, while the sparse component focuses on salient pixels, enabling efficient processing of large-scale images.
+To integrate features, a Sparse Feature Prediction Block (SFB) was designed to reconstruct sparse components, and a Sparse Transformer Blocks (STB) [6] was introduced to capture the long-range contextual information of an image.
+In the restoration module, we design a the Physical-Aware Clear Feature Prediction Block (PFB), guided by the atmospheric scattering model [76].
+The PFB derives illumination and reflectance features from images to predict the distribution of light transmission, utilizing learned convolutional sparse coding (LCSC) [46].
+It removes the need for weather-specific model design, enabling simultaneous training of multiple teacher networks in a unified restoration framework.
+Subsequently, distillation learning transfers knowledge from multiple teacher networks to a single student network, enabling the model to handle adverse weather degradations.
+(3) Construction of Large-Scale All-Weather Benchmark: Furthermore, we introduce a large-scale benchmark, AWMM-100k, specifically designed for multi-modality image fusion in adverse weather conditions, aiming to advance research in this challenging field.
 Our dataset consists of 100,000 multi-modality image pairs, comprehensively covering both synthetic and real-world scenes affected by three major weather degradations: rain, haze, and snow.
 To ensure robust evaluation, the data spans multiple degradation levels for each weather type, capturing the diversity and complexity of natural environments.
-This benchmark addresses a critical gap in the field by providing a reliable, diverse, and extensive resource for developing and evaluating fusion methods under real-world adverse weather conditions.
-The contributions of this study are summarized as follows.
-We propose an end-to-end all-weather MMIF framework, effectively fusing and recovering the perturbed multi-modality information.
-Unlike existing MMIF works, to the best of our knowledge, this is the first work to address MMIF and adverse weather removal simultaneously, defining a new challenging task.
-We propose a feature separation-based fusion module that considers the reconstruction of local and global features from sparse and low-rank components.
-We propose a physically-aware clear feature prediction block (PFB) that comprehensively accounts for the restoration of clear features, thereby avoiding a time-consuming empirical network design.
-Leveraging both existing MMIF datasets and our collected data, we built a dataset (AWMM-100k) dedicated to MMIF research under adverse weather.
-It containing 100,000 aligned infrared and visible image pairs with labels for object detection, establishing it as a versatile benchmark for a broad spectrum of vision tasks.
-II Related Work
-II-A DL-based MMIF Methods
+This benchmark addresses a critical gap in the field by providing a diverse, and extensive resource for developing and evaluating fusion methods under real-world adverse weather.
+The significant contributions of this work are summarized below:
+We propose an end-to-end unified all-weather MMIF framework.
+To the best of our knowledge, this is the first work to address MMIF and adverse weather removal simultaneously, defining a new challenging task.
+We propose a feature separation-based fusion module that considers the reconstruction of local and global features from sparse and low-rank aspects, and develop a physically-aware feature prediction block that comprehensively accounts for the restoration of clear features.
+We construct an all-weather MMIF benchmark (AWMM-100k) contains rain, haze, and snow scenes with different level.
+AWMM-100k covers real-world scenes, and has totally 100,000 registered multi-modality image pairs with object detection labels.
+Extensive experiments on adverse weather image fusion, as well as downstream tasks cover depth estimation, semantic segmentation, and object detection, verify the effectiveness and superiority of the proposed method.
+2 Related Work
+2.1 Image Fusion in Conventional Conditions
 DL-based IVIF algorithms [72, 44, 12, 13, 20, 36, 52, 27] demonstrate robust fusion performance in conventional scenes, employing methodologies such as GAN or AE to effectively integrate salient feature information from each modality.
-With the widespread utilization of IVIF technology, researchers increasingly integrate this technology into downstream vision tasks [3, 32, 33, 34, 35].
+For example, Zhao et al.
+[74] introduced a self-supervised framework incorporating equivariant priors to address the lack of ground truth in multimodal image fusion.
+Similarly, Cheng et al.
+[7] exploited underlying visual tasks in digital photography fusion as supervisory signals and proposed a generalized fusion network that enhances cross-task interaction.
+With the widespread utilization of IVIF technology, researchers increasingly integrate this technology into downstream vision tasks [16, 25, 50, 39].
 Their aim is to enhance the extraction of multi-modality features, augmenting semantic information and leveraging fused images to optimize the performance of tasks such as object detection, semantic segmentation, and salient object detection.
-Furthermore, there are many joint multi-tasking algorithms designed to cope with real-world applications [36, 37, 38].
-In practical applications, images captured by cameras may be slightly misaligned and distorted; some researchers have designed correction networks for fusion specifically for such situations [37].
+For instance, Tang et al.
+[50] argued that most image fusion methods overlook the demands of high-level visual tasks, leading to a lack of semantic information in the fused results.
+To address this, they proposed a semantic-aware real-time image fusion network.
+Liu et al.
+[37] noted that existing approaches often treat fusion and downstream tasks as separate processes, which limits the benefits of fusion for subsequent tasks.
+To overcome this, they introduced a discriminative cross-dimensional evolutionary learning framework for IVIF, employing an evolutionary algorithm to optimize the loss function parameters.
+2.2 Image Fusion in Complex Conditions
+Additionally, there are many joint multi-tasking algorithms designed to cope with real-world applications [21, 64, 56].
+In practical applications, images captured by cameras may be slightly misaligned and distorted; some researchers [14] have designed correction networks for fusion specifically for such situations.
 Xu et al.
 [64] designed an IVIF model that could fuse unaligned image pairs, effectively reducing the cost of adopting IVIF technology.
-In some cases, the fixed input and output resolutions of images can limit the development of IVIF-based methods; Li et al.
-[39] designed an IVIF framework specifically for this problem.
+In some cases, fixed input and output resolutions constrain the development of IVIF methods.
+To address this limitation, Li et al.
+[19] To address this limitation, Li et al.
+proposed a meta-learning-based IVIF framework, where a meta-amplification module dynamically predicts amplification filter weights using scale factors as input.
 In addition, some algorithms apply the IVIF task to degraded environments, such as noise [29], low light [49], and low contrast scenes [68].
+For example, Li et al.
+[31] proposed a unified IVIF framework for complex scenes based on optical-transmission decomposition.
+The framework leverages the prior that degradation factors cause scattering during light transmission, allowing the transmission map to capture the distribution of different degradations.
+By separating the degraded features from the source image, the loss of detail in the feature extraction process is avoided.
+Zhang et al.
+[70] combined multi-modal large models to develop a robust, language-driven image fusion algorithm.
+It leverages the denoising mechanism of latent diffusion models to progressively remove degraded information and can locate targets of interest based on language instructions, highlighting specific scene details.
 However, current research does not effectively address the generalization of IVIF algorithms, whether they are single-task or multi-task, to adverse weather conditions.
-III Proposed Method
-III-A Overview
-As shown in the Fig.
-1, AWFusion comprises two main parts: image fusion and image restoration.
-It is important to emphasize that the proposed algorithm is trained in an end-to-end manner.
-First, the multi-modality data affected by adverse weather is sent to the image fusion module for feature encoding.
-To effectively distinguish private and shared features from different modalities and thereby allow the fusion network to overcome the limitations in using a single encoder, a LLRR block is introduced.
-This block decomposes the source image into a low-rank component, which captures global structural information with high pixel correlation, and a sparse component, which represents high-frequency information with faster pixel variations.
-This approach enables the targeted handling of specific modal features and enhances fusion outcomes.
-In the second part, the multi-modality interaction features obtained from the fusion module are fed into the restoration module to infer the original clear scene.
-As shown in Fig.
-1, the restoration module in the model uses a U-Net backbone.
-In the backbone network, we stack five sparse transformer blocks (STB) to extract useful features distributed in adverse weather.
-STB mainly consist of top-k sparse attention (TKSA) and mixed-scale feed-forward network (MSFN).
-Regarding TKSA, it performs feature aggregation to keep the features in an image sparse, and MSFN extracts multiscale local information from the image [20].
-The proposed PFB focuses on refining the spatial information, which is embedded into both sides of the backbone network and enforced before being inputted into subsequent modules, which helps recover the detailed information.
-Finally, the reconstructed fusion results are output through the restoration module.
-Finally, drawing inspiration from Che et al.
-[43], we introduced a distillation learning approach in which distinct teacher models are initially trained for individual weather-related issues.
-Subsequently, knowledge is collaboratively transferred from these multiple teacher models to student models to comprehensively address weather-related challenges.
-III-B Image Fusion Module
-Through the decomposition provided by the LLRR block, our network focuses solely on the fusion of low-rank and sparse features.
+3 Proposed Method
+3.1 Problem Formulation
+Given an infrared and visible image pair captured under adverse weather, our objective is to generate a fused image that retains complementary modality information while recovering a clean visual scene.
+The task can be formulated as:
+where denotes the proposed AWFusion network and represents the learnable parameters.
+This formulation involves two key goals: (1) effective cross-modal feature fusion and (2) restoration of weather-degraded structures and textures.
+Fusion module. To disentangle complementary information, the input images are first decomposed into low-rank and sparse components using the LLRR block:
+where denotes the processing performed by the Learned Low-Rank Representation Blocks, represent the global structure (low-rank components), and contain modality-specific salient details (sparse components).
+The two components are then fused separately to better preserve shared structures and unique characteristics:
+Where and denote the processing performed by the Sparse Transformer Blocks and the Sparse Feature Prediction Blocks, respectively, producing low-rank fused features and sparse fused features , which are then aggregated and fed into the restoration module for clear-scene reconstruction.
+Restoration module.
+The restoration module employs a U-Net backbone augmented with five STB, each consisting of a Top- Sparse Attention (TKSA) mechanism and a Mixed-Scale Feedforward Network (MSFN) [6].
+TKSA aggregates the most informative features while maintaining sparsity, and MSFN extracts multi-scale local information.
+Additionally, the PFB refines spatial information and is embedded on both sides of the backbone to improve feature representation.
+The restored fusion result is then output as the final prediction .
+Finally, to enhance the robustness of model under adverse weather conditions, we introduce a distillation learning strategy [5].
+Specifically, multiple teacher networks are first trained for distinct weather degradations, and their knowledge is then transferred to a single student network to comprehensively address adverse weather effects.
+3.2 Image Fusion Module
+Due to significant differences between modalities and potential semantic conflicts in certain cases, directly fusing them in the spatial domain can lead to insufficient feature interaction, contrast loss caused by modality misalignment, or misalignment artifacts.
+Therefore, decomposing image features prior to fusion facilitates the network in capturing interactions among different types of features, enabling a better understanding of the semantic information.
+Therefore, we decompose the image into low-rank and sparse components through the LLRR blocks.
 Low-rank components reflect the global information and main structure of the image, constrained by certain rank features.
@@
 In contrast, sparse components represent local details such as edges, textures, and high-frequency information.
-To effectively manage these local fusions, we design the Sparse Feature Block (SFB).
-III-B1 SFB
-Sparse feature fusion is performed to effectively extract and integrate local features specific to an infrared image () and a visible image ().
-The fused feature map consists of two parts: the sparse feature for and the sparse feature for .
-is modeled as follows:
-where and are the sparse filters of related to and , respectively, and denotes the number of filters. Subsequently, Eq. 1 can be transformed into the following optimization problem:
+To effectively manage these local fusions, we design the SFB.
+3.2.1 Sparse Feature Prediction Block (SFB)
+Sparse feature fusion is performed to effectively extract and integrate local features specific to infrared images () and visible images ().
+Here, we employ an optimization principle inspired by LCSC to guarantee that the learned sparse filters represent meaningful high-frequency image details.
+This is motivated by the physical observation that edges and textures dominate local image variations, and proper modeling is necessary to preserve them in the fused output.
+The fused feature map consists of two parts: the sparse feature for and the sparse feature for , modeled as:
+where and are the convolutional filters for and , respectively.
+To derive a tractable optimization problem from the convolutional representation in Eq.
+4, we first express the convolutions in matrix-vector form.
+Each filter and is converted into a Toeplitz matrix and , while the corresponding feature maps and are vectorized.
+This allows the fused feature to be equivalently written as a linear combination of sparse components:
+Next, to ensure that the sparse features accurately reconstruct the original images, we introduce reconstruction fidelity terms for each modality, measured by the -norm of the difference between the source images and their reconstructions:
+Finally, to encourage sparsity in the learned features and suppress redundant activations, we add an regularization term for each sparse feature.
+Combining the fidelity and sparsity terms naturally leads to a joint optimization problem that can be solved using convolutional sparse coding.
+The sparse feature responses are then obtained by solving:
 Similar to [10], and are solved as follows:
@@
 2.
-III-C Image Restoration Module
+3.3 Image Restoration Module
 The second phase aims to achieve image restoration in the presence of three types of adverse weather conditions (rain, snow, and haze).
-Based on the atmospheric scattering model, adverse weather conditions such as rain, haze, and snow induce scattering and absorption of light by particles in the atmosphere.
-Consequently, these atmospheric phenomena lead to comparable image degradation across a spectrum of unfavorable conditions [44].
+Based on the Atmospheric Scattering Model (ASM), adverse weather conditions such as rain, haze, and snow induce scattering and absorption of light by particles in the atmosphere.
+Consequently, these atmospheric phenomena lead to comparable image degradation across a spectrum of unfavorable conditions [77, 31].
 For instance, during heavy rainfall, the accumulation of distant rain lines and atmospheric water particles creates a ”rain curtain” effect.
@@
 The above conclusion provide a research basis to realize deweathering in the same framework.
-Therefore, for the restoration module, an atmospheric scattering model is combined to simulate the process of image degradation, and a PFB is proposed.
-III-C1 PFB
+To leverage this unified representation, we construct the PFB guided by the ASM.
+Specifically, the network is designed to predict the key parameters of the ASM, including the atmospheric light and the transmission map, using learnable neural modules.
+By estimating these parameters, the network can effectively model the scattering and absorption effects induced by adverse weather.
+Subsequently, the clear image features are reconstructed by combining the predicted atmospheric light and transmission map, enabling the recovery of scene information in a physically meaningful and interpretable manner.
+This approach enables the network to learn from data while incorporating physical degradation priors.
+3.3.1 Physical-Aware Clear Feature Prediction Block (PFB)
 The construction of the PFB was guided by the physical mechanism of image degradation and transformation of the atmospheric scattering model.
@@
 where denotes the convolution operator, and represents the Hadamard product.
-Subsequently, by introducing the matrix-vector forms of , , , , and , which are denoted as K, J, I, A, and D, respectively, Eq.
+Subsequently, by introducing the matrix-vector forms of , , , , and , which are denoted as , , , , and , respectively, Eq.
 10 can be redefined as follows:
@@
 Additionally, atmospheric particles can cause light to be reflected, refracted, and absorbed, impacting the overall illumination of the scene.
-Consequently, for the prediction of the feature , by analyzing changes in reflectance and illumination of objects in an image, we can infer changes in light transmission, thus realizing estimation of the image transmission information.
-We disassemble into a combination of reflectance and illumination features, which is modeled as follows:
-where and are the reflectance and illumination filters of , respectively. Subsequently, we compute the reflectance and illumination feature responses by solving the following optimization problem:
+Consequently, for the prediction of the feature , by analyzing changes in reflectance and illumination of objects in an image, we can infer changes in light transmission, thus realizing estimation of the transmission information.
+We disassemble into reflectance and illumination components, modeled as:
+where and are the convolutional filters for reflectance and illumination, respectively.
+To derive a tractable optimization formulation from the convolutional decomposition in Eq.
+13, we first express the convolutions in matrix-vector form.
+Each filter and is converted into a Toeplitz matrix and , while the corresponding feature maps and are vectorized.
+This allows the decomposition to be equivalently written as a linear combination of reflectance and illumination components:
+Then, we introduce reconstruction fidelity terms for both components, measured by the -norm of the difference between and their reconstructions:
+Lastly, to encourage sparsity in the learned reflectance and illumination features and suppress redundant activations, we add an regularization term for each feature map.
+Combining the fidelity and sparsity terms leads to a joint optimization problem that can be solved using convolutional sparse coding.
+Accordingly, the reflectance and illumination feature responses are obtained by solving:
+Furthermore, We emphasize that our objective is adverse weather restoration rather than strict Retinex-based intrinsic decomposition.
+In our framework, reflectance and illumination serve as two feature subspaces for estimating transmission-related cues in the atmospheric scattering model.
+Although a shared sparsity prior is applied, convolutional filters and loss fusion still guide the model to learn distinct responses, effectively separating scene details from low-frequency degradation.
 Reflectance Feature Prediction Block (RFB)
-The goal of RFB is used to compute the reflectance feature. According to [19], we fix to update by rewriting Eq. 9 as:
+The RFB is designed to extract reflectance features, we formalize this as a convolutional sparse coding problem. According to [10], we fix to update by rewriting Eq. 16 as:
 Eq. 17 is a standard convolutional sparse coding problem that can be solved using LCSC [46] by rewriting Eq. 17 as:
 where is the stack of feature responses , is the update of at the -th iteration, and are the learnable convolutional layers related to the filters , and represents the soft thresholding operator.
-Similar to SFB, we expand the iterations in Eq.
+Similarly, we expand the iterations in Eq.
 18 into a network as shown in Fig.
@@
 Illumination Feature Prediction Block (IFB)
-The goal of IFB is to compute the illuminance feature in a hazy image, the structure of this network is similar to that of DFB. By fixing to update , Eq. 9 can be rewritten as:
+The goal of IFB is to compute the illuminance feature in the degraded image, the structure of this network is similar to that of RFB. By fixing to update , Eq. 16 can be rewritten as:
 By using LCSC to solve the above problem, Eq. 19 can be rewritten as follows:
@@
 2.
-III-D Loss Function
-The loss function integrates two components.
+3.4 Loss Function
+The loss function consists of two main components: the first corresponds to the training loss of the teacher networks, while the second pertains to the training loss of the student network.
+The former is primarily based on standard image fusion losses, whereas the latter incorporates knowledge distillation.
+3.4.1 Teacher Networks Loss Function
+This loss function integrates two components.
 The first part focuses on the loss function used to ensure the quality of the fusion results, while the second part addresses the loss function constraining the generation of reflectance and illumination features.
-The fusion loss includes structural similarity index (SSIM), mean square error (MSE) and color consistency loss.
+The fusion loss includes structural similarity index (SSIM) [59], mean square error (MSE) and color consistency loss.
 The calculation for SSIM can be expressed as follows:
@@
 where and represent the height and width of the image, respectively. The MSE loss () is as follow:
-Colour consistency loss is used to ensure that the fused image maintains consistent colours with the .
+Colour consistency loss [68] is used to ensure that the fused image maintains consistent colours with the .
 We transform the image into the YCbCr color space and apply a constraint using the Euclidean distance between the Cb and Cr channels.
@@
 To effectively constrain the generation of the reflectance feature and the illuminance feature in Eq.
-8, we propose a reconstruction loss inspired by retina theory.
+13, we propose a reconstruction loss inspired by retina theory.
 According to this theory, a natural image can be viewed as the product of its reflectance and luminance components.
@@
 The reconstruction loss can be expressed as follows:
-The total loss of AWFusion is denoted as:
+The total loss of teacher network is denoted as:
+3.4.2 Student Networks Loss Function
+The student network learns from teacher networks trained for three types of adverse weather via knowledge distillation.
+Its training objective builds on the teacher network loss while incorporating a Collaborative Knowledge Transfer loss () and a Soft Contrastive Regularization loss ()[5].
+The consists of the Projection Feature Error loss () and Projection Feature Validation losses .
+projects teacher and student feature maps into a shared lower-dimensional space to mitigate the impact of architectural differences on knowledge transfer, while ensures that the feature error computed in is based on informative and valid representations.
+In essence, this loss does not treat teacher features as hard targets; instead, the projection and validation mechanisms address feature space misalignment across networks.
+The treats outputs from well-performing teacher networks as positive (soft) targets and the original degraded images as negative samples, enabling the student network to learn effective weather-specific restoration solutions and facilitating smoother knowledge integration during early training.
+In summary, the total loss of student network is denoted as:
 | Aspect | Description |
@@
 | Task Labels | Object Detection |
-III-E MMIF Benchmark for Adverse Weather
+3.5 MMIF Benchmark for Adverse Weather
+Dataset Overview
 Current MMIF datasets lack comprehensive scene coverage under adverse weather.
@@
 Each type of weather (rain, haze, and snow) is categorized into heavy, medium, and light levels.
-Specific descriptions are shown in Table I.
-The AWMM-100k dataset was collected using a DJI M30T drone, equipped with high-resolution visible and thermal imaging cameras to capture data in complex scenes.
+Specific descriptions are shown in Table 1.
+The AWMM-100k dataset was collected using a DJI M30T drone, equipped with visible and thermal imaging cameras.
+Challenges in Data Collection
 Collecting a large amount of aligned paired multi-modality data and corresponding ground truth in real adverse weather is challenging, leading most current research to rely on synthetic data for training.
@@
 Observations of the data in Fig.
-3 allow us to draw the following conclusions: 1) While the visible image exhibits degradation due to adverse weather conditions, the corresponding infrared image lacks explicit weather features, showcasing only varying degrees of degradation commensurate with the extent of image disturbance, such as diminished contrast and detail.
-2) The depth of the scene affects the extent of light refraction and scattering, resulting in greater blurring in distant scenes compared to closer ones.
-Based on the above conclusions, we propose an adaptive degradation strategy for infrared images based on scene depth.
+3 allow us to draw the following conclusions:
+Visible images show clear weather-related degradation, while infrared images exhibit only general contrast and detail reduction without explicit weather cues.
+The depth of the scene affects the extent of light refraction and scattering, resulting in greater blurring in distant scenes compared to closer ones.
+Building on these observations, we proposed a adaptive degradation strategy for multi-modality images.
+Infrared Degradation Strategy
 First, we transpose the image into the Fourier domain and use the depth map as the weights for degradation.
@@
 In order to suppress too small weights, we add , and to points with pixel values less than according to the three degrees of light, medium and heavy, respectively.
+Visible Image Degradation Strategy
 Degraded visible images generation process is illustrated in Fig.
@@
 We employ a phenomenological pipeline [61] to introduce haze into source images, while rain and snow images are generated using Photoshop software utilizing masks based approach [4].
-In addition, Fig.
-4 presents examples of multi-modality data captured in real-world scenarios under adverse weather.
-Moreover, this benchmark provides labels for object detection task.
-These attributes render it a versatile resource for investigating a broad spectrum of visual tasks.
-IV Experiment
-IV-A Experimental Configurations
+To ensure the realism of synthetic rain and snow images in the AWMM-100K dataset, we extract masks from real-world rain and snow video sequences, which are then used in the rendering process to generate degraded images.
+This approach preserves the natural shape, orientation, motion blur, density, and spatial structure of raindrops and snowflakes, such as linear streaks for rain and scattering or aggregation patterns for snow [4, 26].
+Consequently, the synthetic data exhibit high statistical and visual realism, enabling the model to generalize effectively to real-world scenarios.
+As shown in Fig.
+4, the degradations in the synthetic data of AWMM-100K closely match those observed in real degraded scenes.
+Fig.
+5 presents examples of multi-modality data captured in real-world scenarios.
+Moreover, this benchmark provides labels for object detection.
+These attributes make it a versatile resource for exploring a wide range of visual tasks.
+4 Experiment
+4.1 Experimental Setup
 We performed fusion experiments on the AWMM-100k dataset under adverse weather (rain, haze, and snow).
@@
 This comparison setup allows for a more effective demonstration of the performance of the proposed algorithm.
-IV-B Training Details
+4.2 Training Details
 For the training process, three specialized teacher networks and one student network were trained under different weather conditions.
@@
 The size of the convolutional kernel used in each convolutional layer is nine.
-All experiments were carried out using the PyTorch framework on a system equipped with two NVIDIA 4090 GPUs.
-IV-C Qualitative Comparison
-IV-C1 Comparison on synthetic data
+All the experiments were conducted on a system equipped with two NVIDIA 4090 GPUs.
+Source Method Pub.
+↑ ↑ ↑ ↑ ↓ ↑ ↑ ↑ ↑ U2Fusion[73] TPAMI 2020 0.2548 0.4524 0.7769 0.4565 901.4884 0.2985 0.3631 1.0276 4.4955 LRRNet[23] TPAMI 2023 0.3194 0.5046 0.7003 0.3593 682.5369 0.4250 0.1708 0.7897 5.8903 MURF[64] TPAMI 2023 0.2570 0.3839 0.5732 0.3218 1523.9926 0.2285 0.2699 0.9151 4.5988 TIM[39] TPAMI 2024 0.2057 0.3498 0.4630 0.3844 827.1437 0.3035 0.1585 0.9651 6.1997 CoCoNet[35] IJCV 2023 0.1526 0.2232 0.4021 0.4530 1099.7753 0.2129 0.1756 1.3002 7.6143 TarDal[34] CVPR 2022 0.3563 0.3859 0.5739 0.2710 653.6011 0.3826 0.3053 1.5556 6.2024 IGNet[25] MM 2023 0.3529 0.4921 0.7232 0.4570 919.9543 0.4420 0.3515 1.5791 6.0812 DRMF[47] MM 2024 0.3108 0.4170 0.5844 0.5292 929.2306 0.3555 0.2743 0.8937 6.2728 MUFusion[8] InfFus 2023 0.3438 0.4696 0.7860 0.5102 1039.5903 0.4199 0.3959 1.4061 5.6906 GT CrossFuse[22] InfFus 2024 0.4057 0.5718 0.7132 0.4655 1840.1690 0.4648 0.3832 1.0081 5.9859 AWFusion-T 0.3582 0.6855 0.8444 0.4613 464.7867 0.4839 0.3693 1.6311 6.3160 Rain AWFusion-S - 0.3334 0.6428 0.8335 0.4691 476.3436 0.4618 0.3528 1.6010 6.2777 U2Fusion[73] TPAMI 2020 0.3654 0.4275 0.7236 0.5051 840.3653 0.4086 0.3911 1.1512 5.7205 LRRNet[23] TPAMI 2023 0.4154 0.4775 0.7646 0.4455 463.4354 0.5218 0.3046 0.8130 6.6553 MURF[64] TPAMI 2023 0.2589 0.3324 0.6395 0.4055 1305.4273 0.2277 0.3124 0.9431 5.3601 TIM[39] TPAMI 2024 0.3020 0.3983 0.6499 0.4135 507.1376 0.4078 0.2833 0.8978 7.1141 CoCoNet[35] IJCV 2023 0.2664 0.2446 0.6157 0.4647 556.1065 0.3547 0.2987 1.4288 7.6564 TarDal[34] CVPR 2022 0.4109 0.3832 0.7308 0.3710 337.3376 0.4366 0.4009 1.4207 6.8040 IGNet[25] MM 2023 0.3271 0.4117 0.5225 0.4234 1293.6847 0.4110 0.2683 1.4465 6.2330 DRMF[47] MM 2024 0.5203 0.6138 0.7530 0.4463 477.8087 0.5026 0.4182 1.0414 6.5732 MUFusion[8] InfFus 2023 0.3629 0.4200 0.7236 0.4908 2079.8455 0.4451 0.3723 1.2420 6.5892 GT CrossFuse[22] InfFus 2024 0.4542 0.5423 0.7301 0.4559 2474.6507 0.4847 0.3850 0.9437 6.5648 AWFusion-T 0.4497 0.5889 0.7949 0.4838 374.9215 0.5585 0.4078 1.2209 6.9558 Haze AWFusion-S - 0.4283 0.5471 0.7738 0.4610 462.8585 0.5322 0.3901 1.1719 6.9469 U2Fusion[73] TPAMI 2020 0.3035 0.4330 0.7349 0.4547 1031.8395 0.3606 0.3666 1.0589 5.1796 LRRNet[23] TPAMI 2023 0.3574 0.4651 0.7154 0.3737 661.1762 0.4719 0.2361 0.8301 6.1878 MURF[64] TPAMI 2023 0.2440 0.3547 0.5581 0.3482 1613.7692 0.2180 0.2676 0.8499 5.0984 TIM[39] TPAMI 2024 0.2433 0.3626 0.5181 0.3739 820.5891 0.3465 0.2105 0.8977 6.5951 CoCoNet[35] IJCV 2023 0.2420 0.2580 0.5238 0.4512 828.9278 0.3277 0.2593 1.4450 7.5726 TarDal[34] CVPR 2022 0.3511 0.3718 0.6238 0.2903 553.2230 0.4171 0.3453 1.4684 6.4487 IGNet[25] MM 2023 0.3268 0.4577 0.5938 0.4324 1141.8494 0.4356 0.2986 1.5142 6.0336 DRMF[47] MM 2024 0.3948 0.4684 0.6747 0.5055 729.0171 0.4649 0.3417 1.2065 6.2139 MUFusion[8] InfFus 2023 0.3543 0.4336 0.7342 0.4953 1699.4374 0.4372 0.3787 1.2984 6.1981 GT CrossFuse[22] InfFus 2024 0.4090 0.5187 0.6995 0.4567 2421.3634 0.4507 0.3724 0.9919 6.1354 AWFusion-T 0.3580 0.5591 0.8179 0.4827 246.8077 0.4985 0.3780 1.5700 6.6173 Snow AWFusion-S - 0.3552 0.5535 0.8119 0.4798 270.1595 0.4937 0.3734 1.4937 6.6531
+Source Method Pub.
+Background Car Person Bike Curve Car Stop Guardrail Color cone Bump mIoU U2Fusion[73] TPAMI 2020 98.39 92.88 73.25 73.03 52.82 73.75 86.2 40.31 83.2 74.87 LRRNet[23] TPAMI 2023 98.51 93.18 69.62 73.89 65.18 77.95 84.59 35.95 88.96 76.43 MURF[64] TPAMI 2023 99.13 95.56 83.54 86.65 76.27 88.21 93.48 76.25 90.59 87.74 TIM[39] TPAMI 2024 98.22 91.17 67.94 69.4 59.53 67.74 83.42 30.59 85.59 72.62 CoCoNet[35] IJCV 2023 98.3 92.03 69.31 70.66 57.72 73.62 87.49 29.44 84.9 73.72 TarDal[34] CVPR 2022 98.39 92.29 70 68.47 62.07 74.59 87.9 35.65 85.42 75.01 IGNet[25] MM 2023 99.12 95.61 83.09 87.22 74.97 87.96 93.1 76.13 90.9 87.6 DRMF[47] MM 2024 99.16 95.14 82.36 88.2 84.27 90.68 91.67 80.35 92.08 89.32 MUFusion[8] InfFus 2023 99.23 95.17 83.84 88.43 84.1 88.9 92.94 80.7 86.12 88.82 GT CrossFuse[22] InfFus 2024 98.04 90.3 67.43 68.51 58.58 71.81 68.13 29.92 68.7 69.04 Rain 99.32 96.49 84.77 88.76 86.35 90.45 93.91 78.45 91.82 90.04 Haze 99.19 95.69 81.51 81.94 85.21 89.5 91.57 72.73 93.57 87.88 Snow AWFusion-S - 99.25 95.92 81.24 86.4 86.77 90.36 93.74 78.04 92.32 89.34
+Source Method Pub.
+People Car Bus Lamp Motorcycle Truck mAP@0.5 mAP@[0.5:0.95] U2Fusion[73] TPAMI 2020 0.816 0.903 0.901 0.716 0.711 0.801 0.808 0.512 LRRNet[23] TPAMI 2023 0.785 0.909 0.912 0.764 0.721 0.808 0.817 0.484 MURF[64] TPAMI 2023 0.815 0.885 0.881 0.711 0.675 0.8 0.794 0.498 TIM[39] TPAMI 2024 0.731 0.898 0.878 0.657 0.656 0.769 0.765 0.48 CoCoNet[35] IJCV 2023 0.791 0.889 0.889 0.696 0.641 0.762 0.778 0.49 TarDal[34] CVPR 2022 0.792 0.877 0.857 0.585 0.62 0.713 0.741 0.462 IGNet[25] MM 2023 0.793 0.859 0.836 0.532 0.554 0.721 0.716 0.439 DRMF[47] MM 2024 0.699 0.897 0.896 0.787 0.697 0.784 0.793 0.494 MUFusion[8] InfFus 2023 0.803 0.906 0.908 0.734 0.694 0.802 0.808 0.507 CrossFuse[22] InfFus 2024 0.692 0.891 0.871 0.671 0.674 0.771 0.762 0.473 GT AWFusion-S - 0.791 0.922 0.93 0.804 0.726 0.821 0.832 0.534
+4.3 Qualitative Comparison
+4.3.1 Comparison on synthetic data
 Figs.
-5, 6 and 7 present the fusion results of different algorithms on the synthetic dataset.
+6, 7 and 8 present the fusion results of different algorithms on the synthetic dataset.
 Notably, only the proposed algorithm performs fusion under adverse weather conditions, while the comparison methods are fused in ideal scenes.
@@
 As shown in Figs.
-5, 6 and 7, all comparison methods achieve good fusion results in ideal scenes.
-Four methods—U2Fusion, LRRNet, MURF, and IGNet—produce fusion results with low contrast.
+6, 7 and 8, all comparison methods achieve good fusion results in ideal scenes.
+Four methods, including U2Fusion, LRRNet, MURF, and IGNet, produce fusion results with low contrast.
 In contrast, CoCoNet and MUFusion enhance the source image features, leading to some redundancy and loss of detail in the thermal radiation information in the fusion results.
@@
 However, in nighttime scenes, the TarDal method exhibits a loss of detail.
-Additionally, both MUFusion and CrossFuse show poor color retention of the sky, introducing excessive infrared information into the region, which is not compatible with the human visual system.
+Additionally, both MUFusion and CrossFuse show poor color retention of the sky, introducing excessive infrared information, which is not compatible with the human visual system.
 Under severe weather interference, the challenge not only arises from weather-related artifacts on the visible images but also from the low contrast of infrared images and the loss of thermal radiation information.
@@
 However, this does not affect the semantic content of the overall image, and the student model remains close to the teacher model in terms of overall contrast.
-IV-C2 Comparison on real-world data
+To demonstrate the capability of proposed method in multi-modal feature interaction and extraction, we conduct a comparison under degradation-free conditions.
+As shown in Fig.
+9, all methods receive identical inputs.
+The results indicate that our method still maintains superior fusion performance and does not fail when applied to clean scenes, despite being trained on images affected by adverse weather.
+4.3.2 Comparison on real-world data
 Fig.
-9, we present the fusion results of all methods in a real haze scene.
+11, we present the fusion results of all methods in a real haze scene.
 Additionally, we evaluate the fusion performance of different methods using two downstream tasks: object detection and depth estimation.
@@
 However, as seen in Fig.
-8, all dehaze methods resulted in varying degrees of color distortion, suggesting that feeding such distorted inputs into the different fusion methods could lead to unfair comparisons.
-Therefore, we provided all fusion methods with uniform input images.
+10, all dehaze methods resulted in varying degrees of color distortion, suggesting that feeding such distorted inputs into the different fusion methods could lead to unfair comparisons.
+Therefore, no restoration algorithm was incorporated into the fusion method to remove interfering factors from the source images.
 As shown in Fig.
-9, in the first set of fusion comparison tests, CoCoNet, TarDal, and MUFusion effectively retain the significant thermal radiation information from the infrared map.
+11, in the first set of fusion comparison tests, CoCoNet, TarDal, and MUFusion effectively retain the significant thermal radiation information from the infrared map.
 These three methods, which tend to enhance infrared data to achieve higher contrast, can address some challenges in real haze scenes.
@@
 While all methods preserved depth in the foreground, many, such as U2Fusion, LRRNet, CoCoNet, and MUFusion, struggled with accurate depth estimation for distant pedestrians obscured by haze.
-Additionally, haze affected the performance of TIM, IGNet, and DRMF, as these methods failed to maintain contrast between the roadside fence and its surroundings, leading to edge blurring in the depth map.
-In summary, the proposed algorithm outperforms current SOTA methods in both fusion performance and semantic information retention when addressing real-world scene challenges.
+In summary, the proposed algorithm outperforms current SOTA methods in both fusion performance and semantic information retention when addressing real-world scene challenges, and demonstrates excellent performance in downstream tasks.
 It demonstrates excellent performance in downstream tasks, such as object detection and depth estimation, indicating its high practical application value in real-world.
-Source Method Pub.
-↑ ↑ ↑ ↑ ↓ ↑ ↑ ↑ ↑ U2Fusion[2] TPAMI 2020 0.2548 0.4524 0.7769 0.4565 901.4884 0.2985 0.3631 1.0276 4.4955 LRRNet[4] TPAMI 2023 0.3194 0.5046 0.7003 0.3593 682.5369 0.4250 0.1708 0.7897 5.8903 MURF[38] TPAMI 2023 0.2570 0.3839 0.5732 0.3218 1523.9926 0.2285 0.2699 0.9151 4.5988 TIM[35] TPAMI 2024 0.2057 0.3498 0.4630 0.3844 827.1437 0.3035 0.1585 0.9651 6.1997 CoCoNet[50] IJCV 2023 0.1526 0.2232 0.4021 0.4530 1099.7753 0.2129 0.1756 1.3002 7.6143 TarDal[3] CVPR 2022 0.3563 0.3859 0.5739 0.2710 653.6011 0.3826 0.3053 1.5556 6.2024 IGNet[32] MM 2023 0.3529 0.4921 0.7232 0.4570 919.9543 0.4420 0.3515 1.5791 6.0812 DRMF[51] MM 2024 0.3108 0.4170 0.5844 0.5292 929.2306 0.3555 0.2743 0.8937 6.2728 MUFusion[52] InfFus 2023 0.3438 0.4696 0.7860 0.5102 1039.5903 0.4199 0.3959 1.4061 5.6906 GT CrossFuse[53] InfFus 2024 0.4057 0.5718 0.7132 0.4655 1840.1690 0.4648 0.3832 1.0081 5.9859 AWFusion-T 0.3582 0.6855 0.8444 0.4613 464.7867 0.4839 0.3693 1.6311 6.3160 Rain AWFusion-S - 0.3334 0.6428 0.8335 0.4691 476.3436 0.4618 0.3528 1.6010 6.2777 U2Fusion[2] TPAMI 2020 0.3654 0.4275 0.7236 0.5051 840.3653 0.4086 0.3911 1.1512 5.7205 LRRNet[4] TPAMI 2023 0.4154 0.4775 0.7646 0.4455 463.4354 0.5218 0.3046 0.8130 6.6553 MURF[38] TPAMI 2023 0.2589 0.3324 0.6395 0.4055 1305.4273 0.2277 0.3124 0.9431 5.3601 TIM[35] TPAMI 2024 0.3020 0.3983 0.6499 0.4135 507.1376 0.4078 0.2833 0.8978 7.1141 CoCoNet[50] IJCV 2023 0.2664 0.2446 0.6157 0.4647 556.1065 0.3547 0.2987 1.4288 7.6564 TarDal[3] CVPR 2022 0.4109 0.3832 0.7308 0.3710 337.3376 0.4366 0.4009 1.4207 6.8040 IGNet[32] MM 2023 0.3271 0.4117 0.5225 0.4234 1293.6847 0.4110 0.2683 1.4465 6.2330 DRMF[51] MM 2024 0.5203 0.6138 0.7530 0.4463 477.8087 0.5026 0.4182 1.0414 6.5732 MUFusion[52] InfFus 2023 0.3629 0.4200 0.7236 0.4908 2079.8455 0.4451 0.3723 1.2420 6.5892 GT CrossFuse[53] InfFus 2024 0.4542 0.5423 0.7301 0.4559 2474.6507 0.4847 0.3850 0.9437 6.5648 AWFusion-T 0.4497 0.5889 0.7949 0.4838 374.9215 0.5585 0.4078 1.2209 6.9558 Haze AWFusion-S - 0.4283 0.5471 0.7738 0.4610 462.8585 0.5322 0.3901 1.1719 6.9469 U2Fusion[2] TPAMI 2020 0.3035 0.4330 0.7349 0.4547 1031.8395 0.3606 0.3666 1.0589 5.1796 LRRNet[4] TPAMI 2023 0.3574 0.4651 0.7154 0.3737 661.1762 0.4719 0.2361 0.8301 6.1878 MURF[38] TPAMI 2023 0.2440 0.3547 0.5581 0.3482 1613.7692 0.2180 0.2676 0.8499 5.0984 TIM[35] TPAMI 2024 0.2433 0.3626 0.5181 0.3739 820.5891 0.3465 0.2105 0.8977 6.5951 CoCoNet[50] IJCV 2023 0.2420 0.2580 0.5238 0.4512 828.9278 0.3277 0.2593 1.4450 7.5726 TarDal[3] CVPR 2022 0.3511 0.3718 0.6238 0.2903 553.2230 0.4171 0.3453 1.4684 6.4487 IGNet[32] MM 2023 0.3268 0.4577 0.5938 0.4324 1141.8494 0.4356 0.2986 1.5142 6.0336 DRMF[51] MM 2024 0.3948 0.4684 0.6747 0.5055 729.0171 0.4649 0.3417 1.2065 6.2139 MUFusion[52] InfFus 2023 0.3543 0.4336 0.7342 0.4953 1699.4374 0.4372 0.3787 1.2984 6.1981 GT CrossFuse[53] InfFus 2024 0.4090 0.5187 0.6995 0.4567 2421.3634 0.4507 0.3724 0.9919 6.1354 AWFusion-T 0.3580 0.5591 0.8179 0.4827 246.8077 0.4985 0.3780 1.5700 6.6173 Snow AWFusion-S - 0.3552 0.5535 0.8119 0.4798 270.1595 0.4937 0.3734 1.4937 6.6531
-IV-D Quantitative Comparison
-IV-D1 Fusion Performance Comparison
+Methods Pub.
+Restor.
+↑ ↑ ↑ ↑ ↓ ↑ ↑ ↑ ↑ Rain Text-DiFuse [69] NIPS 24 0.2442 0.3645 0.5283 0.3909 777.6507 0.1962 1.2938 0.3329 6.9609 ControlFusion [48] NIPS 25 0.3507 0.4750 0.6901 0.4361 739.1370 0.2996 1.1971 0.4576 7.0261 GIFNet [7] CVPR 25 0.2975 0.3686 0.8040 0.4834 672.4981 0.3474 1.3440 0.3544 5.5197 SAGE [60] CVPR 25 0.2286 0.5086 0.7878 0.3964 465.1593 0.2734 1.2986 0.3670 5.4524 OmniFuse [70] TPAMI 25 AdaIR [9] 0.2191 0.4027 0.5863 0.3774 549.3715 0.1771 1.0588 0.3340 6.9013 AWFusion-T 0.3580 0.6855 0.8444 0.4613 464.7867 0.3693 1.6311 0.4839 6.3160 AWFusion-S _ _ 0.3334 0.6428 0.8335 0.4691 476.3436 0.3528 1.6010 0.4618 6.2777 Haze Text-DiFuse [69] NIPS 24 0.3137 0.3463 0.6622 0.4391 597.2099 0.3035 1.2850 0.4021 7.3050 ControlFusion [48] NIPS 25 0.3841 0.4032 0.7309 0.4652 547.6077 0.3331 1.1490 0.4767 6.9131 GIFNet [7] CVPR 25 0.3124 0.3181 0.6878 0.4784 847.9477 0.3463 1.2118 0.3478 6.2780 SAGE [60] CVPR 25 0.3416 0.3948 0.7347 0.4553 557.8396 0.3316 1.1922 0.4336 6.6047 OmniFuse [70] TPAMI 25 AdaIR [9] 0.2638 0.3448 0.6483 0.4226 514.9563 0.2558 1.0287 0.3446 7.0612 AWFusion-T 0.4497 0.5889 0.7949 0.4838 374.9215 0.4078 1.2209 0.5585 6.9558 AWFusion-S _ _ 0.4283 0.5471 0.7738 0.4610 462.8585 0.3901 1.1719 0.5322 6.9469 Snow Text-DiFuse [69] NIPS 24 0.3047 0.3633 0.6263 0.4264 502.4444 0.2821 1.5002 0.4130 7.0752 ControlFusion [48] NIPS 25 0.4184 0.4610 0.7487 0.4855 427.3205 0.3554 1.3418 0.5194 7.0112 GIFNet [7] CVPR 25 0.3282 0.3305 0.7706 0.4854 482.2309 0.3722 1.3885 0.3882 6.0665 SAGE [60] CVPR 25 0.3149 0.4752 0.7910 0.4421 275.8248 0.3474 1.4294 0.4585 6.0712 OmniFuse [70] TPAMI 25 AdaIR [9] 0.2729 0.3845 0.6254 0.4114 318.5334 0.2515 1.1877 0.3977 7.0768 AWFusion-T 0.3580 0.5591 0.8179 0.4827 246.8077 0.3780 1.5700 0.4985 6.6173 AWFusion-S _ _ 0.3552 0.5535 0.8119 0.4798 270.1595 0.3734 1.4937 0.4937 6.6531
+Methods Pub.
+Restor.
+↑ ↑ ↑ ↑ ↓ ↑ ↑ ↑ ↑ Rain Text-DiFuse [69] NIPS 24 0.2485 0.3622 0.5280 0.4076 770.4863 0.2000 1.2776 0.3384 6.9881 ControlFusion [48] NIPS 25 0.3809 0.4677 0.6769 0.4420 740.8956 0.2956 1.1693 0.4531 7.0698 GIFNet [7] CVPR 25 0.3075 0.3747 0.8062 0.4804 670.4288 0.3512 1.3463 0.3650 5.5507 SAGE [60] CVPR 25 0.2355 0.5356 0.7925 0.4072 465.1666 0.2872 1.3023 0.3793 5.5068 OmniFuse [70] TPAMI 25 MODEM [57] 0.2248 0.4068 0.5903 0.3912 545.6923 0.1845 1.0391 0.3434 6.9319 AWFusion-T 0.3580 0.6855 0.8444 0.4613 464.7867 0.3693 1.6311 0.4839 6.3160 AWFusion-S _ _ 0.3334 0.6428 0.8335 0.4691 476.3436 0.3528 1.6010 0.4618 6.2777 Haze Text-DiFuse [69] NIPS 24 0.3513 0.3752 0.7024 0.4728 516.7434 0.3380 1.3969 0.4577 7.3805 ControlFusion [48] NIPS 25 0.4402 0.4761 0.7560 0.4981 553.4258 0.3673 1.2045 0.5511 7.1785 GIFNet [7] CVPR 25 0.3530 0.3366 0.7344 0.5007 1090.9966 0.3834 1.2959 0.4085 6.3176 SAGE [60] CVPR 25 0.4026 0.5124 0.7906 0.4826 626.5782 0.3935 1.3033 0.5216 6.6250 OmniFuse [70] TPAMI 25 MODEM [57] 0.2866 0.3705 0.6982 0.4621 529.6896 0.2834 1.0791 0.4028 7.1742 AWFusion-T 0.4497 0.5889 0.7949 0.4838 374.9215 0.4078 1.2209 0.5585 6.9558 AWFusion-S _ _ 0.4283 0.5471 0.7738 0.4610 462.8585 0.3901 1.1719 0.5322 6.9469 Snow Text-DiFuse [69] NIPS 24 0.3089 0.3672 0.6289 0.4312 492.1872 0.2848 1.4871 0.4207 7.0754 ControlFusion [48] NIPS 25 0.4150 0.4619 0.7486 0.4816 422.5371 0.3538 1.3377 0.5220 7.0211 GIFNet [7] CVPR 25 0.3227 0.3350 0.7693 0.4821 487.3009 0.3665 1.3761 0.3874 6.0588 SAGE [60] CVPR 25 0.3172 0.4913 0.7937 0.4453 262.5657 0.3486 1.4287 0.4653 6.0729 OmniFuse [70] TPAMI 25 MODEM [57] 0.2763 0.3902 0.6289 0.4182 311.3442 0.2540 1.1914 0.4052 7.0759 AWFusion-T 0.3580 0.5591 0.8179 0.4827 246.8077 0.3780 1.5700 0.4985 6.6173 AWFusion-S _ _ 0.3552 0.5535 0.8119 0.4798 270.1595 0.3734 1.4937 0.4937 6.6531
+4.4 Quantitative Comparison
+4.4.1 Fusion Performance Comparison
 We selected 50 pairs of infrared and visible source images from the AWMM 100k dataset for each of the three types of adverse weather: rain, haze, and snow, as the test set.
 The ground truth corresponding to the degraded images was used as input for all comparison methods.
-Table II presents the results of all methods across nine objective metrics, with the first column indicating the source image types used by each method.
+Table 2 presents the results of all methods across nine objective metrics, with the first column indicating the source image types used by each method.
 The top three scores for each metric are highlighted.
-Notably, Table II shows that the proposed method, even when performing fusion under adverse weather conditions, outperforms comparison methods operating under ideal conditions.
-In rain, haze, and snow, the proposed teacher model ranks in the top three for eight or more metrics, while the student model ranks in the top three for , , and metrics, respectively.
+Notably, Table 2 shows that the proposed method, even when performing fusion under adverse weather conditions, outperforms comparison methods operating under ideal conditions.
+In rain, haze, and snow, the proposed teacher model ranked in the top three across more than eight evaluation metrics.
+Meanwhile, the student model achieved top-three rankings on seven, four, and seven metrics under the respective conditions.
 By specifically analyzing the scores of each metric, we can draw the following conclusions:
@@
 As a result, snow is slightly easier to process than rain and haze.
-The proposed algorithm does not rank among the top three in the metric for haze scenes.
-This metric evaluates image quality based on texture, local structural changes, and noise sensitivity, suggesting that the performance of algorithm in extracting local texture under haze is somewhat weaker than in other conditions.
-(3) For the metrics and , which are inspired by human perception, the proposed method performs exceptionally well.
-This aligns with the results shown in Figs.
-5, 6 and 7 and Fig.
-9, where the method effectively integrates multi-modality information while eliminating interference, ensuring that the fusion results are not distorted by adverse weather or enhanced in ways that conflict with human visual perception.
-In summary, the performance of the proposed algorithm in quantitative comparisons further validates its practical value in handling severe weather conditions, demonstrating excellent fusion capabilities.
-Source | Method | Pub. | Background | Car | Person | Bike | Curve | Car Stop | Cuardrail | Color cone | Bump | mIoU |
-U2Fusion[2] | TPAMI 2020 | 98.39 | 92.88 | 73.25 | 73.03 | 52.82 | 73.75 | 86.2 | 40.31 | 83.2 | 74.87 | |
-LRRNet[4] | TPAMI 2023 | 98.51 | 93.18 | 69.62 | 73.89 | 65.18 | 77.95 | 84.59 | 35.95 | 88.96 | 76.43 | |
-MURF[38] | TPAMI 2023 | 99.13 | 95.56 | 83.54 | 86.65 | 76.27 | 88.21 | 93.48 | 76.25 | 90.59 | 87.74 | |
-TIM[35] | TPAMI 2024 | 98.22 | 91.17 | 67.94 | 69.4 | 59.53 | 67.74 | 83.42 | 30.59 | 85.59 | 72.62 | |
-CoCoNet[50] | IJCV 2023 | 98.3 | 92.03 | 69.31 | 70.66 | 57.72 | 73.62 | 87.49 | 29.44 | 84.9 | 73.72 | |
-TarDal[3] | CVPR 2022 | 98.39 | 92.29 | 70 | 68.47 | 62.07 | 74.59 | 87.9 | 35.65 | 85.42 | 75.01 | |
-IGNet[32] | MM 2023 | 99.12 | 95.61 | 83.09 | 87.22 | 74.97 | 87.96 | 93.1 | 76.13 | 90.9 | 87.6 | |
-DRMF[51] | MM 2024 | 99.16 | 95.14 | 82.36 | 88.2 | 84.27 | 90.68 | 91.67 | 80.35 | 92.08 | 89.32 | |
-MUFusion[52] | InfFus 2023 | 99.23 | 95.17 | 83.84 | 88.43 | 84.1 | 88.9 | 92.94 | 80.7 | 86.12 | 88.82 | |
-GT | CrossFuse[53] | InfFus 2024 | 98.04 | 90.3 | 67.43 | 68.51 | 58.58 | 71.81 | 68.13 | 29.92 | 68.7 | 69.04 |
-Rain | 99.32 | 96.49 | 84.77 | 88.76 | 86.35 | 90.45 | 93.91 | 78.45 | 91.82 | 90.04 | ||
-Haze | 99.19 | 95.69 | 81.51 | 81.94 | 85.21 | 89.5 | 91.57 | 72.73 | 93.57 | 87.88 | ||
-Snow | AWFusion-S | - | 99.25 | 95.92 | 81.24 | 86.4 | 86.77 | 90.36 | 93.74 | 78.04 | 92.32 | 89.34 |
-Source | Method | Pub. | People | Car | Bus | Lamp | Motorcycle | Truck | mAP@0.5 | mAP@[0.5:0.95] |
-U2Fusion[2] | TPAMI 2020 | 0.816 | 0.903 | 0.901 | 0.716 | 0.711 | 0.801 | 0.808 | 0.512 | |
-LRRNet[4] | TPAMI 2023 | 0.785 | 0.909 | 0.912 | 0.764 | 0.721 | 0.808 | 0.817 | 0.484 | |
-MURF[38] | TPAMI 2023 | 0.815 | 0.885 | 0.881 | 0.711 | 0.675 | 0.8 | 0.794 | 0.498 | |
-TIM[35] | TPAMI 2024 | 0.731 | 0.898 | 0.878 | 0.657 | 0.656 | 0.769 | 0.765 | 0.48 | |
-CoCoNet[50] | IJCV 2023 | 0.791 | 0.889 | 0.889 | 0.696 | 0.641 | 0.762 | 0.778 | 0.49 | |
-TarDal[3] | CVPR 2022 | 0.792 | 0.877 | 0.857 | 0.585 | 0.62 | 0.713 | 0.741 | 0.462 | |
-IGNet[32] | MM 2023 | 0.793 | 0.859 | 0.836 | 0.532 | 0.554 | 0.721 | 0.716 | 0.439 | |
-DRMF[51] | MM 2024 | 0.699 | 0.897 | 0.896 | 0.787 | 0.697 | 0.784 | 0.793 | 0.494 | |
-MUFusion[52] | InfFus 2023 | 0.803 | 0.906 | 0.908 | 0.734 | 0.694 | 0.802 | 0.808 | 0.507 | |
-CrossFuse[53] | InfFus 2024 | 0.692 | 0.891 | 0.871 | 0.671 | 0.674 | 0.771 | 0.762 | 0.473 | |
-GT | AWFusion-S | - | 0.791 | 0.922 | 0.93 | 0.804 | 0.726 | 0.821 | 0.832 | 0.534 |
-IV-D2 Semantic Segmentation Performance Comparison
+The proposed algorithm is not among the top three in the metric for haze scenes.
+This metric evaluates image quality based on texture and local structural changes, suggesting that the performance of algorithm in extracting local texture under haze is somewhat weaker than in other conditions.
+4.4.2 Semantic Segmentation Comparison
 The proposed all-weather MMIF algorithms are designed to support real-world applications, such as autonomous driving, surveillance, and UAV navigation.
@@
 To comprehensively evaluate the fusion performance under three types of severe weather, we introduced a downstream task.
-In Table III, we compare the performance of various algorithms using BANet [61] as the segmentation network.
+In Table 3, we compare the performance of various algorithms using BANet [41] as the segmentation network.
 Unlike comparison methods, which receive clean source images as input, the proposed algorithm performs fusion under severe weather conditions.
-As shown in Table III, the proposed algorithm ranks first in mean intersection-over-union (mIoU) for rain scenes and second for snow scenes.
+As shown in Table 3, the proposed algorithm ranks first in mean intersection-over-union (mIoU) for rain scenes and second for snow scenes.
 Although the mIoU score for the haze scenario is not in the top three, it still secures fifth place.
 These experiments demonstrate that the proposed algorithm retains high-level semantic information in scenes with interference and accurately recognizes key targets such as roads, pedestrians, and vehicles.
-IV-D3 Object Detection Performance Comparison
+Methods Pub.
+Restor.
+↑ ↑ ↑ ↑ ↓ ↑ ↑ ↑ ↑ Snow and Haze ControlFusion [48] NIPS 25 MODEM [57] 0.3408 0.3230 0.7097 0.4499 383.7061 0.3097 1.0466 0.4581 7.1471 GIFNet [7] CVPR 25 0.2950 0.2811 0.6826 0.4579 717.2248 0.3304 1.2148 0.3595 6.6378 SAGE [60] CVPR 25 0.3212 0.3630 0.7247 0.4445 443.1403 0.3259 1.1365 0.4524 7.0594 AWFusion-S _ _ 0.3578 0.3590 0.7494 0.4584 519.5943 0.3698 1.3985 0.4656 7.1554 Rain and Haze ControlFusion [48] NIPS 25 MODEM [57] 0.3385 0.3453 0.7036 0.4350 385.4859 0.2995 0.9922 0.4680 7.3067 GIFNet [7] CVPR 25 0.3023 0.2956 0.6924 0.4529 662.5418 0.3292 1.2082 0.3685 6.6517 SAGE [60] CVPR 25 0.3182 0.3880 0.7221 0.4296 436.8800 0.3138 1.0881 0.4516 7.0836 AWFusion-S _ _ 0.3775 0.3999 0.7530 0.4454 549.1318 0.3722 1.4212 0.4919 7.1886 Rain and Snow ControlFusion [48] NIPS 25 MODEM [57] 0.3551 0.3203 0.7258 0.4594 537.3284 0.3155 1.2346 0.4869 7.0407 GIFNet [7] CVPR 25 0.2924 0.2826 0.6588 0.4622 1121.0535 0.3113 1.1707 0.3739 6.6209 SAGE [60] CVPR 25 0.3138 0.3781 0.7579 0.4419 628.5957 0.3124 1.2184 0.4821 6.7914 AWFusion-S _ _ 0.3382 0.3622 0.6955 0.4699 258.0324 0.3198 1.3459 0.4910 7.1263
+4.4.3 Object Detection Performance Comparison
 In real-world applications, systems must adapt to varying weather and environmental conditions.
@@
 This demonstrates not only its effectiveness in harsh weather but also its strong generalization ability in clean settings.
-In Table IV, we evaluated the object detection accuracies of all methods on the M3FD dataset using YOLOv7 [10].
-As shown in Table IV, the proposed algorithm achieves the highest detection accuracy across all five categories, with both mAP@0.5 and mAP@[0.5:0.95] also obtaining the top scores.
+In Table 4, we evaluated the object detection accuracies of all methods on the M3FD dataset using YOLOv7 [55].
+As shown in Table 4, the proposed algorithm achieves the highest detection accuracy across all five categories, with both mAP@0.5 and mAP@[0.5:0.95] also obtaining the top scores.
 These results confirm that the proposed algorithm retains sufficient information and detail for accurate localization and identification in object detection.
 Moreover, it enhances detection accuracy by preserving and leveraging complementary information.
-IV-E Extension to Multiple Adverse Weather Disturbances
+4.5 Extension to Comparisons with “Restoration + Fusion” pipelines
+To further validate the effectiveness of our method, we extend our experiments to comparisons with “Restoration + Fusion” pipelines.
+Specifically, we first apply image restoration algorithms to recover the degraded visible images, and then feed the restored visible images together with clean infrared images into SOTA IVIF methods.
+In contrast, our model directly takes degraded multi-modal images as input.
+We use two restoration methods, AdaIR [9] and MODEM [57], and compare our method with five advanced IVIF algorithms: Text-DiFuse [69], ControlFusion [48], GIFNet [7], SAGE [60], and OmniFuse [70].
+Qualitative results are shown in Fig.
+12, while quantitative results are reported in Table 5 and Table 6.
+For fairness, both restoration algorithms were retrained on the AWMM-100K dataset.
+From Fig.
+12, we can observe that although restoration algorithms can remove most degradation, the remaining artifacts and detail loss are inevitably passed to the subsequent fusion stage.
+In rainy scenes, the fusion results of Text-DiFuse, ControlFusion, and OmniFuse show varying degrees of detail loss, leading to insufficient scene contrast.
+GIFNet carries noticeable artifacts produced during restoration, which may interfere with downstream tasks such as target recognition and scene understanding.
+In hazy scenes, similar issues persist.
+GIFNet attempts to compensate for the lost details by overly enhancing infrared information, but this causes the final image to deviate from human visual perception.
+Since Text-DiFuse and OmniFuse cannot perceive degradation cues, they amplify the residual degradation from the restoration stage, further reducing image quality.
+In snowy scenes, snowflakes not only occlude structural details but also introduce effects similar to overexposure.
+As a result, GIFNet and SAGE fail to preserve scene brightness, leading to incomplete depth information.
+In contrast, the proposed method consistently delivers robust fusion performance across all three adverse conditions.
+It preserves fine details, maintains color fidelity, and effectively removes degradation, outperforming all “Restoration + Fusion” pipelines.
+Table 5 and Table 6 present the quantitative comparisons using AdaIR-restored and MODEM-restored visible images.
+As reflected in the qualitative analysis, Text-DiFuse and OmniFuse obtain relatively low scores due to their significant loss of details.
+The proposed method surpasses all competing approaches across the three weather conditions and achieves the highest scores on most evaluation metrics.
+Overall, the proposed method outperforms the “Restoration + Fusion” pipelines in both qualitative and quantitative comparisons, further demonstrating its advantages for real-world applications.
+4.6 Extension to Multiple Weather Disturbances
 Compound degradation increases scene complexity, as rain, haze, and snow exhibit distinct physical characteristics: rain and snow cause localized occlusion, while haze reduces overall contrast.
@@
 To evaluate the generalization ability of the proposed algorithm under severe weather conditions, we conducted extended experiments with composite degradation.
-We used DT [62] as a preprocessing model to remove interference from visible images.
-In the comparison method, fusion is performed on the preprocessed image, while the proposed algorithm directly processes the fusion result on the original source image.
+In the comparison methods, the fusion operation is performed on pre-processed images.
+We adopt MODEM, a state-of-the-art adverse-weather removal model, to eliminate degradations in the visible images before fusion.
+Since MODEM is trained on AWMM-100K, it is capable of identifying and removing all three types of degradations.
+In contrast, our method directly generates fused results from the original degraded multi-modal inputs, without relying on any pre-processing.
 As shown in Fig.
-10, when two disturbances coexist in the source image, DT can only address one at a time, despite its effectiveness handle individual interferences.
+13, when two types of degradations appear simultaneously in the source images, the degradation removal capability of MODEM is significantly reduced.
 The residual degradation is carried over to the fusion result, leading to suboptimal outcomes.
-In contrast, the proposed method reconstructs high-fidelity scene details under complex weather conditions, preserving key features while avoiding issues like insufficient brightness or information loss.
-IV-F Ablation Experiment
-We performed an ablation analysis on the six modules to verify their individual roles and effectiveness.
+In contrast, the proposed method reconstructs high-fidelity scene details under complex weather conditions, preserving key features while avoiding issues like information loss.
+Table 7 presents the quantitative results of different methods under multiple weather degradations.
+As shown, the proposed method achieves the highest scores on six or more metrics across all three types of composite degradation.
+4.7 Ablation Experiment
+We performed an ablation analysis on the six modules to verify their effectiveness.
 The modules were grouped into four experimental sets, with the results visualized in Fig.
-11 and quantitatively compared in Table V.
-IV-F1 Effectiveness Analysis of the Fusion Module
+14 and quantitatively compared in Table 8.
+4.7.1 Effectiveness Analysis of the Fusion Module
 The image fusion module plays a critical role in the end-to-end framework by enabling multi-modality feature interaction.
@@
 As illustrated in the first row of Fig.
-11, removing the fusion module disrupts this interaction, causing salient infrared features to be occluded by blurred visible details, as seen in the green zoomed-in area.
+14, removing the fusion module disrupts this interaction, causing salient infrared features to be occluded by blurred visible details, as seen in the green zoomed-in area.
 Additionally, the absence of feature interaction introduces low-contrast IR information into the fusion output, making it difficult to reconstruct color information and resulting in slight color distortion, highlighted in the red zoomed-in area.
-IV-F2 Effectiveness Analysis of the PFB
-The PFB module, built on the atmospheric scattering model, effectively mitigates weather-induced interference and reconstructs clear scene features.
-We performed an ablation analysis.
-As shown in the second row of Fig.
-11, removing the PFB module leaves residual degradation in the fusion result.
-In texture-rich regions, such as leaves, the network struggles to distinguish between degradation artifacts and actual textures, as highlighted in the red zoomed-in area, leading to detail loss and blurred edges.
 | w/o PFB | 0.5439 | 0.7917 | 579.3986 | 0.3048 | 6.2609 |
@@
 | AWFusion | 0.6855 | 0.8444 | 464.7867 | 0.3993 | 6.3160 |
-IV-F3 Effectiveness Analysis of the IFB and RFB
+4.7.2 Effectiveness Analysis of the PFB
+The PFB module, built on the atmospheric scattering model, effectively mitigates weather-induced interference and reconstructs clear scene features.
+We performed an ablation analysis.
+As shown in the second row of Fig.
+13, removing the PFB module leaves residual degradation in the fusion result.
+In texture-rich regions, such as leaves, the network struggles to distinguish between degradation artifacts and actual textures, as highlighted in the red zoomed-in area, leading to detail loss and blurred edges.
+4.7.3 Effectiveness Analysis of the IFB and RFB
 To improve the ability of network to estimate light transmission, we designed the IFB and RFB modules based on retina theory and constrained feature generation by reconstructing the loss.
 As shown in the third row of Fig.
-11, removing these modules results in reduced lighting in the fusion output, likely due to inaccurate transmission map estimation and insufficient luminance compensation.
+13, removing these modules results in reduced lighting in the fusion output, likely due to inaccurate transmission map estimation and insufficient luminance compensation.
 This makes the absence of light recovery more noticeable.
-Additionally, Fig.
-4 shows a significant drop in the , , and metrics.
-IV-F4 Effectiveness Analysis of the Fusion Strategies
+Additionally, Table 8 shows a significant drop in the , , and metrics.
+4.7.4 Effectiveness Analysis of the Fusion Strategies
 To leverage the low-rank and sparse features decomposed by the LLRR block, we designed the SFB to fuse sparse components and the STB to integrate low-rank components through context extraction.
@@
 As more feature channels are extracted from the source image, the simple fusion strategy impedes effective cross-modal interaction, restricting the extraction of salient and complementary features.
-This causes pixel redundancy in the fused output, reflected in the excessively high metric in Table V.
+This causes pixel redundancy in the fused output, reflected in the excessively high metric in Table 8.
 The enlarged region in Fig.
-11 further highlights increased noise in ground, complicating object recognition.
-V Conclusion
-In this study, we propose an all-weather MMIF method, focusing on the relationship between multi-modality image fusion and image restoration.
-Our framework shifts the network’s objective from pixel-perfect recovery to maximizing the representation of key scene information.
-The end-to-end design comprises two main modules: a fusion module and a restoration module.
-The fusion module decouples features from multiple modalities into low-rank and sparse components, which are merged interactively using specialized fusion strategies.
-Guided by a physical model, the restoration module restores clear details from blur pixels, ensuring high-quality fusion results.
+13 further highlights the increased noise in the ground, which complicates object recognition.
+5 Discussion
+Limitations and Future Work: The proposed algorithm offers an effective solution for IVIF under adverse weather conditions.
+While it demonstrates strong fusion performance, our experiments reveal that its computational efficiency leaves room for further improvement.
+Specifically, for an input of size 256×256, the algorithm requires 193 ms of inference time, with 889 GFLOPs and 19.359 million parameters.
+In future work, we aim to optimize the computational efficiency of the model, thereby enhancing its applicability in real-time and resource-constrained scenarios.
+Conclusion: In this study, we propose an all-weather MMIF method, focusing on the relationship between multi-modality image fusion and image restoration.
+Our framework shifts the goal of the network from pixel-perfect recovery to maximizing the representation of key information from the scene.
+The end-to-end design comprises two main modules: a fusion module that decouples multi-modal features into low-rank and sparse components for interactive merging, and a restoration module guided by physical model to restore clear details from blur pixels, ensuring high-quality results.
 Additionally, we introduce a large-scale benchmark dataset, AWMM-100k, tailored for multi-modality image processing under adverse weather conditions.
 Experimental results confirm that the AWFusion surpasses existing methods in both adverse and normal conditions, with outstanding performance in downstream tasks, demonstrating its strong practical value.
+Acknowledgement
+This research was supported by the National Natural Science Foundation of China (No.
+62201149), the Natural Science Foundation of Guangdong Province (No.
+2024A1515011880), the Basic and Applied Basic Research of Guangdong Province (No.
+2023A1515140077), the Research Fund of Guangdong-HongKong-Macao Joint Laboratory for Intelligent Micro-Nano Optoelectronic Technology (No.
+2020B1212030010).